### v1

# Personalized Generative Low-light Image Denoising and Enhancement

Abstract
While smartphone cameras today can produce astonishingly good photos, their performance in low light is still not completely satisfactory because of the fundamental limits in photon shot noise and sensor read noise. Generative image restoration methods have demonstrated promising results compared to traditional methods, but they suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low. Recognizing the availability of personalized photo galleries on users’ smartphones, we propose Personalized Generative Denoising (PGD) by building a diffusion model customized for different users. Our core innovation is an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery. This ID-consistent physical buffer provides a strong prior that can be integrated with the diffusion model to restore the degraded images, without the need of fine-tuning. Over a wide range of low-light testing scenarios, we show that PGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches.
Figure 1: Using gallery photos from a user’s smartphone, we can restore low-light, noisy facial images. Our method produces finer details and better identity compared to current state-of-the-art face restoration approaches like DR2 [53] and GFP-GAN [50].
22footnotetext: Project page: https://genai-restore.github.io/DiffPGD/.
1 Introduction
With the astonishing development of smartphones, it is safe to say that smartphone cameras today have surpassed digital single-lens reflex (DSLR) cameras in both popularity and functionality. However, the small form factor of smartphone cameras puts a tight constraint on the aperture size, hence limiting the amount of light a CMOS pixel can detect. In a low-light imaging environment, this creates a fundamental limitation to how short the minimum exposure needs to be and how much signal-to-noise ratio (SNR) the sensor can support. While some of the noise seen at low light can be mitigated using better CMOS technology (e.g., correlated double sampling to reduce the read noise [40] and deep-well pump gate to reduce the stray capacitance between the floating diffusion and the transfer gate [39, 38]), the Poisson / Bose-Einstein statistics due to random photon arrivals is a problem created by mother nature that cannot be solved by even the ideal sensors. Therefore, image denoising and enhancement by exploiting the internal structures of images become a necessary task for all smartphone cameras.
From a signal-processing perspective, degradation caused by low light can be roughly approximated via
$\mathbf{y}=\text{ADC}\big{(}\text{Poisson}(\alpha\mathbf{x}+\eta)\big{)}+\epsilon,$ (1)
where $\mathbf{x}$ is the unknown clean image, $\alpha$ is the sensor gain, $\eta$ is the dark current, and $\epsilon$ is the read noise. The inverse problem associated with low-light denoising is to recover $\mathbf{x}$ from $\mathbf{y}$. Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [35], total variation [43, 3], wavelets [8] to non-local means [2], BM3D [9], and recently convolution neural networks [24, 61, 48, 5, 1, 17, 55], transformers [34, 37, 60], and diffusion models [25, 58, 66]. A common theme across existing methods is that they all focus on a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images. While they perform well in moderately difficult problems, the generality of these methods cannot be extended to heavily corrupted images. Particularly in the context of human face recovery, the restored results often lack real identity and exhibit artifacts. This is attributed to the ill-posed nature of the problem and the lack of appropriate constraints.
Why Gallery Photos? Smartphone cameras today store hundreds if not thousands of a user’s photos, captured at different times, in different places, and under different lighting conditions. While these images have many variations, they are all about the same person(s). Therefore, if the imaging goal is to take a photo of this user, the gallery on the phone would be the best source to build a prior $p(\mathbf{x})$. The situation is summarized in Fig. 2. In the context of diffusion-based image restoration, the original solution space can be large because many candidate solutions are consistent with the noisy observation. The gallery provides a strong constraint to the search problem. This allows us to search for better quality images with high precision of the person’s identity.
Figure 2: The restoration of inputs degraded by noise and low-light conditions is highly ill-posed. By incorporating additional high-quality gallery photos of the same identity, we significantly reduce the solution space, thereby achieving improved identity consistency in the restored images.
There are, however, two technical questions:
  • What class of images to focus on: In this paper, we will focus on human faces because of their relevance to most smartphone users although our algorithm can be equally applied to other image classes. We assume that the gallery photos have been selected and are reasonably informative. Product-level engineering such as pre-processing and selection of the gallery photos is important but beyond the scope of this paper.
  • How to incorporate this class-specific $p(\mathbf{x})$ into the restoration model: The bigger (and harder) question is that given the gallery photos, how do we efficiently extract the prior information and improve the restoration? We assume that the gallery set is not too big and so we cannot train a restoration model from scratch. We want to avoid fine-tuning as much as possible during inference. Moreover, we want to make sure that the person’s identity is preserved.
Physical Buffers to the Rescue. Given the gallery photos, what kind of prior information would be useful for restoration? Advancements in computer vision have made it possible to extract detailed facial physical buffers—including albedo and normal maps—from a person’s gallery of images[42, 63]. These physical buffers capture intrinsic properties such as skin color and surface geometry, effectively encoding an individual’s unique identity and fine facial details. (See Fig. 3.) At the same time, it also eliminates the influence of environmental lighting, pose, and other identity-independent variables. We will leverage this rich prior information to improve restoration.
To convince readers that physical buffers could be useful to our denoising task, we show in Fig. 3 a toy experiment. We consider three types of degradation: low-resolution (LR), motion blur, and noise. In each case, we run an existing physical buffer extraction method DECA[14] to extract the albedo and surface normal of the input. We observe that despite the different types of degradation, the physical buffers (albedo and surface normal) are fairly accurate. The only failure case is when the noise is extremely strong. Therefore, if we can develop solutions to manage the heavy noise case, we will be able to use physical buffers as a signal prior to guiding the denoising process.
Contributions. Our contributions are threefold:
  1. 1.
    Personalized Low-Light Denoise and Enhancement Framework: We introduce a new framework that leverages diffusion models and person-specific priors from photo galleries to address identity loss in restored low-light facial images.
  2. 2.
    Physical Identity Buffers: We propose the use of physical identity buffers from existing face galleries to enrich the restoration process, enabling accurate reconstruction of human faces under severe degradation.
  3. 3.
    Easy Deployment: Our method operates easily without requiring fine-tuning for individual users, making it practical for real-world applications and scalable to multiple users.
Extensive experiments demonstrate that our approach outperforms existing techniques in both visual quality and identity preservation, providing a significant advancement in low-light facial image restoration.
2 Related Works
Low-light Image Enhancement and Denoising. Low-light images often suffer from significant noise due to low signal-to-noise ratios (SNR), in addition to reduced brightness and contrast. Traditional low-light denoising techniques often focus on exploiting image structures, ranging from early methods like Wiener filtering [35], total variation [43, 3], and wavelets [8] to non-local means [2] and BM3D [9]. More recently, deep learning approaches, including convolutional neural networks [24, 61, 48, 5, 1, 17, 55, 65], transformers [34, 37, 60], and diffusion-based models [25, 58, 66, 41], have emerged as promising solutions for low-light denoising. Low-light image enhancement (LLIE) aims to improve the brightness and contrast of images captured in poorly lit conditions. Traditional approaches [15, 30, 28] often leverage heuristic or physical models to enhance image quality, while learning-based methods [27, 54, 23, 4, 16, 36, 51, 57, 21] have demonstrated greater effectiveness when supported by large-scale datasets. Recent advancements [44, 37, 41, 56] in this field have increasingly focused on jointly addressing denoising and enhancement tasks, aiming to improve image quality while effectively suppressing noise to maintain natural appearance. However, to the best of our knowledge, no specific approach currently exists for joint enhancement and denoising of facial images in low-light conditions.
Figure 3: For images affected by mild low resolution, motion blur, or noise, DECA [14] can accurately extract physical buffers. However, DECA fails to obtain precise physical buffers for images degraded by severe low light and noise.
Deep Face Restoration. CNN/Transformer-based face restoration methods [52, 13, 22], such as RestoreFormer [52], typically rely on paired low- and high-quality image datasets to learn resolution enhancement and noise reduction. While effective, these methods often struggle under severe degradation and lack the capability to synthesize realistic details in complex scenarios.
In recent years, generative priors have been increasingly utilized in face restoration to better address real-world degradations. GAN-based methods, such as GFP-GAN [50], leverage StyleGAN2 [47] to enhance facial details, while diffusion-based models [53, 6, 12, 46, 49] have further expanded restoration capabilities. For instance, BFRDiffusion [6] uses Stable Diffusion to enhance low-quality images by adding high-fidelity details, and Ding et al. [12] employ a pretrained diffusion model that denoises degraded inputs while preserving identity by fine-tuning on a person’s gallery photos. MGBFR [46] combines text prompts and reference images with a dual-control adapter to retain accurate facial attributes and identity. Although these generative methods are effective, they often rely on multi-stage training or finetunning, which can limit the efficient use in some real-world applications.
Reference Prior for Face Restoration. In face restoration, reference images [46, 12, 31, 32] provide essential priors for recovering detailed and high-frequency facial features. These priors guide the model to generate realistic restorations. Reference-based methods can be divided into two types: single reference and gallery-based images. The single reference approach [46] uses one image to guide the restoration, focusing on enhancing facial details from a specific reference. In contrast, the gallery-based approach [12] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration.
3 Method
3.1 Restoration by Diffusion
A diffusion model approximates the clean image distribution $p_{\theta}(x_{0})$ by learning a model $\theta$ that reverses the noise-adding process. In Denoising Diffusion Probabilistic Models (DDPMs) [20], Gaussian noise is progressively added to a clean image $x_{0}$:
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon,\quad\text{where}\quad\epsilon\sim\mathcal{N}(0,I)$ (2)
The reverse process then removes the noise to reconstruct the clean image. After training, for any given time $t$ and the corresponding noisy image $x_{t}$, the model can iteratively denoise by sampling from $p(x_{0}|x_{t})$.
In restoration tasks, like low-light denoising, the goal is to recover a high-quality image $x_{0}$ from a low-quality input $y_{0}$. Unlike pure generation, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $y_{0}$. During training, the model uses $y_{0}$ as a condition, and once the model is trained, for any given time $t$ and the corresponding noisy image $x_{t}$, it can iteratively denoise by sampling from $p(x_{0}|x_{t},y_{0})$ using the trained model.
Figure 4: Path to the idea. Using facial physical buffers and gallery images can effectively preserve human face identity information.
Figure 5: The overall architecture of the proposed method. Our core idea is to use ID-consistent physical buffers, extracted from gallery photos, to constrain the generative space in the diffusion model restoration process. For a high-quality gallery, we use LAP [64] to extract the albedo and normal information for each photo and apply adaptive aggregation to fuse the entire gallery. The extracted albedo represents base skin color and texture details, such as skin tone, freckles, and pigmentation, while normal captures facial geometry, including wrinkles, pores, and fine surface variations. In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
3.2 Why Gallery + Physical Buffer?
We formalize our core approach of extracting identity-consistent physical buffers from gallery photos to guide the diffusion model for restoration. This method is based on the observation that using facial physical buffers or gallery images can effectively preserve identity information during restoration tasks involving diffusion models. To evaluate the capability of physical buffers and gallery images in preserving identity information without relying on prior identity knowledge from degraded inputs, we train the following models using a blind scheme, where the degraded input ($y_{0}$) is not used as a condition during training. As illustrated in Fig. 4, the models are designed as follows:
  • Model 1: Trained exclusively on generic images as a purely generative model.
  • Model 2: Trained on generic images while conditioned on the facial physical buffer extracted from the input image. This model incorporates face identity and facial features through the physical buffer.
  • Model 3: Fine-tuned on Model 1 with the user’s gallery photos. This model learns identity features from the target person’s gallery photos during fine-tuning.
  • Model 4: Fine-tuned on Model 2 with the user’s gallery photos. This model integrates facial identity and features from the physical buffer of the input image, as well as the target person’s gallery photos.
We test these models using a simple low-light noisy input. Model 1 generates random faces, as expected for a purely generative model. In contrast, Model 2 produces images with facial features (e.g., nose, mouth, skin color) resembling those of the user. This is because the physical buffers extracted from the degraded user image (refer to Fig. 3) were used as conditioning. These results demonstrate that facial features and identity information can be incorporated by leveraging physical buffers. The outputs from Model 3 are all different user faces, compared to Model 1, showing that fine-tuning with gallery photos can effectively transfer strong identity information. The outputs from Model 4 are the best-expected results, demonstrating the effectiveness of using gallery photos and physical buffers to maintain the person’s identity.
However, despite the powerful combination of gallery images and physical buffer, there are limitations:
  1. (i)
    Fine-tuning: Extracting identity information from gallery photos via fine-tuning is time-consuming and inconvenient. Each time when we switch to a different user, the model must be fine-tuned again. Additionally, since the size of a gallery is usually much smaller than a pre-trained model, fine-tuning can easily overfit.
  2. (ii)
    Extracting physical buffers from the degraded input image: If the input is heavily degraded, the physical buffers extracted from it can be wrong (refer to Fig. 3), and result in terrible results as shown in Fig. 4. Besides, the facial feature extracted from a single image is limited, not to mention from a degraded image.
In conclusion, we want to avoid fine-tuning the model on the gallery and we do not want to extract incorrect physical buffers due to noise.
3.3 ID-Consistent Physical Buffer
To address these limitations, we propose extracting physical buffers directly from existing clean gallery photos and using them as conditions to guide the diffusion model during training. This approach offers significant benefits: clean gallery photos allow us to form comprehensive and robust identity-consistent physical buffers containing detailed facial identity features of the target person. Specifically, We will extract the albedo and normal maps of the target person, which are two key attributes used to describe the physical characteristics of a face.
  • Albedo captures the intrinsic color of an object’s surface—that is, the subject’s color properties without any lighting effects.
  • Normal information is used to describe the fine bumps and details such as skin wrinkles and pores.
By conditioning on these strong identity physical buffers, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction. Consequently, the model is better equipped to preserve the target person’s identity and facial features, even under severe degradation.
3.4 Extracting Physical Buffers
Extracting physical buffers for a person from a set of gallery photos is a more challenging task than extracting these buffers from a single input, as in methods like DECA. This complexity arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses. To ensure these buffers comprehensively represent the person’s identity and facial details, they must be aligned to maintain identity consistency. Fortunately, existing works have addressed this challenge. As shown in the physical buffer extractor module in Fig. 5, we adopt the aggregation net design from the LAP [64] to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
3.5 Model Framework
We name our final proposed model DiffPGD. As illustrated in Fig. 5, the model leverages ID physical buffers (albedo $a^{i}$ and normal $n^{i}$) extracted from person i’s gallery photos as conditions to guide our diffusion model during training directs the generative process toward a personalized space.
Within the aggregation network, it contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting consistent face. To model albedo and normal, two separate aggregation networks denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed. Given a photo gallery of N images $\{\textbf{I}_{k}^{i}\}_{k=1}^{N}$ of person i, each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $x_{k}^{a}$ and $x_{k}^{n}$, respectively. To derive a global representation of the identity for person i based on $\{x_{k}^{a},x_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted, This method learns a weight set $\{w_{k}^{a},w_{k}^{n}\}_{k=1}^{N}$ according to the quality of each image in$\{\textbf{I}_{k}^{i}\}_{k=1}^{N}$. The combined global ID-code $x_{c}^{a}$ and $x_{c}^{n}$ for texture and depth are calculated as:
$x_{c}^{a}=\sum_{k=1}^{N}w_{k}^{a}x_{k}^{a},\quad x_{c}^{n}=\sum_{k=1}^{N}w_{k}^{n}x_{k}^{n}.$ (3)
Then the global ID-code $x_{c}^{a}$ and $x_{c}^{n}$ are then passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $a^{i}$ and normal $n^{i}$.
Additionally, with the low-light noisy image $y_{0}^{i}$ as another condition during training, other features, such as pose, hairstyle, and background inside the image beyond facial features can be effectively captured as well.
After the model is trained, during testing, the process begins with a Gaussian noise image $x_{T}$, which is iteratively transformed into the clean, enhanced image $x_{0}^{i}$ by sampling from $p(x_{0}^{i}|x_{T},y_{0}^{i},a_{i},n_{i})$ using the trained model. Our experiments demonstrate that identity information is effectively preserved with the guidance of these ID-consistent buffers, enabling our model to achieve superior results in human face restoration.
4 Experiments
4.1 Dataset
Training Data. Since no publicly available datasets exist for low-light noisy face images, we simulate such data using the CelebAMask-HQ dataset [29], the InverseISP algorithm [33], and the Poisson-Gaussian noise model Eq. 1. The pipeline involves converting RGB images to their corresponding RAW mosaiced format using InverseISP [33], followed by applying Eq. 1 to simulate low-light noise using realistic camera sensor parameters. Multiple noise levels are generated by varying the light intensity (Lux) to ensure the model is robust across diverse noise conditions. Refer to supplementary for more details on the simulation.
Testing Data. For testing DiffPGD, we utilize both simulated and real-world low-light noisy data. The simulation setup assumes a $12$-bit image sensor with parameters similar to that of Canon DSLR cameras: a Full Well Capacity of $36000\text{e}^{-}$, Quantum Efficiency of $0.42$, dark current of $7\text{e}^{-}$, and read noise of $6\text{e}^{-}$. These parameters ensure realistic noise characteristics for evaluating model performance. We collect the real-world test data using the Sony $\alpha 6400$ DSLR camera at an approximate light intensity of $0.13$ Lux with an ISO of $25600$ and an exposure time of $\frac{1}{8}$ seconds.
4.2 Implementation Details
Our denoising model is built on the ADM architecture [11], and the ID-consistent physical extractor is implemented using the aggregation network from LAP [63]. For training, we use a P2-weighted loss function [7], which measures the discrepancy between the predicted and ground-truth noise. This loss is defined as: $L=\lambda_{t}||\hat{\epsilon}_{t}-\epsilon_{t}||_{2}^{2}$ , where $\lambda_{t}$ serves as a hyperparameter to adjust the loss weight dynamically across different timesteps.
During the training stage, we train our model on the CelebAMask-HQ dataset [29], comprising 30,000 images distributed across 4,516 unique identities(IDs). We use 3318 IDs for training and 200 IDs for simulated testing. Within each ID, there are 3–20 images. The Adam optimizer is applied with a learning rate of 1e-4. Training is performed for 200,000 iterations with a batch size of 32, resulting in a total of 6,400,000 samples seen by the model.
Figure 6: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
Figure 7: Visual comparison on real-captured photo test (zoomed-in area in the input is scaled for visual purposes). Our method outputs the best denoised and enhanced result, and successfully restores fine facial details, such as moles, wrinkles.
4.3 Performance in Simulated Scenarios
In simulated scenarios, we compare our methods with several state-of-the-art works: UTVNet[65], ACENet[56], GFPGAN[50], SNR-aware[57], DR2[53], MIRNet[59] and DiffLL[21]. These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research. Particularly, GFPGAN and DR2 are designed specifically for face image restoration, which is highly competitive with our method. We use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set. We simulated five low-light noisy levels by setting different photons per pixel (PPP), which range from 13 PPP to 65 PPP, the lower the PPP, severe the degradation is.
Quantitative and qualitative results are provided. Firstly, we use PSNR, SSIM, and the perceptual metric (LPIPS[62]) to measure the restored image quality. The results are shown in Table 1, it can be seen that our method has surpassed all previous methods by a large scale. Demonstrating the supreme ability of our proposed DiffPGD model on low light denoising and enhancement problems.
Since our work focuses on face images, we also check on the identity preservation performance. Therefore, besides the PSNR, SSIM, and LPIPS metrics, we adopt the identity score which calculates the cosine similarity of the image features extracted by ArcFace[10] to further evaluate the identity consistency. The results are shown in Table 2. Again, our method has the highest identity score. The advantage of our model on preserving the identity is especially outstanding in heavily degraded cases.
Visualized comparisons are presented in Fig. 6. The image restored from our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistent best. Again, the visual results are consistent with the quantitative metric results, the advantage of our model on preserving the identity is especially outstanding in the heavily degraded cases. The sampled gallery photos of the test subject can be found in Fig. 19. Refer to Sec. B in the supplementary material for more visual results.
  Photons Per Pixel 13 26 39 52 65
[…]
UTVNet[65] 12.08 0.2644 0.7816 13.26 0.3685 0.7169 13.99 0.4332 0.6754 14.50 0.4783 0.6419 14.87 0.5118 0.6148
ACENet[56] 11.25 0.2815 0.7431 12.27 0.4022 0.6725 13.19 0.4670 0.6276 13.87 0.5120 0.5894 14.43 0.5479 0.5577
SNR-aware[57] 10.74 0.1186 0.8193 11.99 0.2031 0.7650 12.34 0.2546 0.7230 12.53 0.2935 0.6852 12.76 0.3271 0.6507
MIRNet[59] 11.64 0.4553 0.7476 13.64 0.5416 0.6491 13.90 0.5724 0.5874 14.12 0.5960 0.5355 14.323 0.6154 0.4934
DiffLL[21] 13.30 0.3472 0.7872 14.66 0.4282 0.7189 15.47 0.4588 0.6736 15.96 0.4805 0.6387 16.25 0.4969 0.6072
DR2[53] 13.50 0.3448 0.7573 15.16 0.5141 0.6430 16.22 0.5655 0.5871 16.85 0.5909 0.5503 17.54 0.6209 0.5181
GFPGAN[50] 13.53 0.4256 0.7049 15.56 0.5446 0.5713 16.60 0.5960 0.5156 17.41 0.6322 0.4748 17.79 0.6563 0.4431
Ours 21.10 0.6287 0.3862 22.80 0.6914 0.3394 23.85 0.7195 0.3156 24.11 0.7354 0.2965 24.38 0.7399 0.2947
Table 1: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset.
  Methods 13 PPP 26 PPP 39 PPP 52 PPP 65 PPP
UTVNet[65] 0.1572 0.3654 0.5201 0.6197 0.6938
ACENet[56] 0.1734 0.4173 0.5769 0.6756 0.7240
SNR-aware[57] 0.1069 0.1921 0.3114 0.4171 0.4971
MIRNet[59] 0.3071 0.5613 0.6694 0.7246 0.7794
DiffLL[21] 0.1751 0.3779 0.5033 0.6107 0.6682
DR2[53] 0.2456 0.5128 0.6358 0.6905 0.7239
GFPGAN[50] 0.2506 0.5392 0.6569 0.7212 0.7615
Ours 0.5995 0.7271 0.7702 0.8044 0.8253
Table 2: Comparisons of identity score with other methods on CelebAMask-HQ dataset.
  Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
UTVNet[65] 12.02 0.3901 0.6954 0.8062
ACENet[56] 11.36 0.3962 0.6817 0.8125
SNR-aware[57] 10.18 0.1922 0.7274 0.6111
MIRNet[59] 12.48 0.4315 0.7132 0.6914
DiffLL[21] 13.18 0.3899 0.7551 0.7686
DR2[53] 14.08 0.1829 0.7350 0.5309
GFPGAN[50] 14.18 0.3714 0.6784 0.7138
Ours 14.02 0.4362 0.5954 0.8248
Table 3: Comparisons with state-of-the-art works on real-captured photo test.
  Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
w/o physical buffers 12.96 0.3801 0.6196 0.7538
w/ physical buffers 14.02 0.4362 0.5954 0.8248
Table 4: Ablation comparisons of using ID physical buffers on real-captured photo test.
4.4 Performance in Real-Life Scenarios
Testing model performance in real-life low-light scenarios is crucial, as real-world conditions often involve uncontrollable factors compared to simulated ones. Since no existing dataset addresses the restoration of real-life low-light noisy face photos, we created a custom test set. Specifically, we printed five photos of five target individuals and captured images of these photos under low-light conditions using a camera with the settings specified in Section 4.1. As shown in Fig. 7, our model demonstrates significantly better recovery of real-captured low-light noisy inputs compared to existing models. In Table 3, we quantitatively compare the performance of our model with other baselines on this real-captured photo test. We observe that all our metrics are better except for GFPGAN’s PSNR. However, when looking at the visual comparison in Fig. 7, we realize that GFPGAN recovers smoother features but loses the true facial details, resulting in a relatively fake-looking face. For further verification, we also photographed real individuals under low-light conditions with the camera. As shown in Fig. 8, DiffPGD outperforms existing methods by a large margin. These results confirm that DiffPGD effectively leverages prior information from a target person’s gallery photos to guide the diffusion model in learning a personalized generative space. During restoration, the model accurately directs the output toward this clean personalized space. The gallery photos of the real-captured male subject are displayed in Fig. 9. Other test subjects’ gallery photos can be found in Fig. 19.
Figure 8: (Left) Environmental setup, real-life low-light noisy photo taken in the dark room. (Right) Visual comparison on the real-world captured SUBJECT.
Figure 9: Male SUBJECT’s gallery (face areas are cropped from the subject’s photos).
5 Ablation Study
ID-Consistent Physical Buffers for Identity Preservation. We have shown that the proposed model DiffPGD exhibits outstanding performance in both simulated and real-world testing scenarios. In Table 4, we compare the quantitative results of training our model with and without ID physical buffers in real-world test cases. The results demonstrate that using physical buffers largely increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
6 Conclusion
In this work, we addressed the challenges of low-light image restoration by introducing Personalized Generative Denoising (PGD), a novel approach that leverages user-specific photo galleries to improve denoising performance. By incorporating an identity-consistent physical buffer, we provided a strong prior that enhances the diffusion model’s ability to restore degraded images while avoiding the need for model fine-tuning. Our experiments demonstrated that PGD outperforms existing diffusion-based denoising methods, delivering superior results across various testing scenarios. This approach not only mitigates false content generation in low-SNR conditions but also bridges the gap in low-light imaging quality, paving the way for personalized and more robust image restoration techniques.
Acknowledgment This research was supported, in part, by a research contract from Samsung Research America, the National Science Foundation (NSF), and in part by the Intelligence Advanced Research Projects Activity (IARPA) under Contract No. 2022-21102100004, and in part by the National Science Foundation under the grants CCSS-2030570 and IIS-2133032. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Supplementary Material
Figure 9: Simulation process for generating low-light noisy face images.
Appendix A Dataset
As mentioned in Section $4.1$ of the main paper, since no open-sourced low-light noisy datasets are available, we adopt a simulation approach to generate low-light noisy face images for training our model. Specifically, we utilize the CelebAMask-HQ dataset’s [29] face images as the ground truth RGB images.
As shown in Fig. 9, the first step in the simulation process involves passing the RGB image through a pre-trained Inverse ISP network [33] to generate a pseudo ground truth RAW mosaic image. For visualization purposes, we display the demosaiced version of this image. We utilize the pretrained weights provided by the authors [33], which were obtained by training the network on iPhone camera images, for unprocessing the RGB images into RAW format.
This pseudo ground truth RAW image is then passed through a Poisson-Gaussian Simulator to simulate the low-light noisy RAW mosaic image. It is noteworthy that during the training process, we vary the camera sensor parameters (in the Poisson-Gaussian simulator) across a wide range of values to enable the model to handle diverse camera sensor noise profiles. The simulation setup assumes a $12$-bit image sensor with parameters similar to those of Canon DSLR cameras. The parameter ranges are summarized in Table 5.
Table 5: Simulation parameters used in the Poisson-Gaussian Simulator.
  Parameter Range
Full Well Capacity $[19000\text{e}^{-},64000\text{e}^{-}]$
Quantum Efficiency $[0.32,0.54]$
Dark Current $[2.2\text{e}^{-},11.7\text{e}^{-}]$
Read Noise $[2.2\text{e}^{-},10.8\text{e}^{-}]$
Photons-per-Pixel (PPP) $[13\,\text{PPP},65\,\text{PPP}]$
These parameters are carefully chosen to ensure realistic noise characteristics during training.
Appendix B Additional Visual Results
In Fig. 16, Fig. 17 and Fig.18, we present additional visual comparisons between our model and baseline methods. Our results consistently demonstrate superior performance, particularly in cases of severe degradation.
Appendix C Results in More Severe Cases
To further evaluate the ability of our model to denoise and enhance extremely degraded photos, we tested it under highly challenging conditions, including both simulated and real-life scenarios.
For simulated cases, We simulate severely degraded test cases with 10 PPP and 5 PPP. As shown in Fig.12 and Fig.13, our model can still perform well in terms of enhancing and denoising the degraded inputs, whereas other models struggled significantly. Besides, due to the extremely low signal level in the input image, we can see some hallucination effect in the hairstyle of the generated outputs. Despite this, thanks to the ID physical buffer extracted from the gallery of the target person, our model could still effectively restore the facial structure of the target person. Quantitative comparisons are presented in Tables 7 and 8.
We also evaluated the model on real-captured photos. Fig. 14 is a visual example of the male subject taken under extremely low light conditions with high noise levels. From the figure, we can see that our proposed model outperforms all baseline methods in both denoising and enhancement. Additionally, the blind-based restoration methods DR2 and GFPGAN completely failed to produce meaningful results due to the severely low signal levels. This supports our observation that blind restoration approaches struggle with images degraded by extremely low light and noise. For these real-captured extreme test cases, we further conducted quantitative evaluations of both image quality and identity preservation, as summarized in Table 6. Since ground truth images are unavailable for real-captured cases, we computed the Fréchet Inception Distance(FID)[19] between the model’s output images and the subject’s gallery photos, which are 4 clean natural photos from the subject’s phone (shown in Fig.9). We calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo. The results indicate that our model consistently achieves superior image quality and identity preservation compared to other methods.
Figure 10: The relationship between the identity (ID) score of the restored images and the number of gallery images used to generate the ID physical buffer.
Figure 11: A network architecture diagram for identity feature restoration using a face encoder.
  Methods FID $\downarrow$ ID Score $\uparrow$
UTVNet[65] 293.2 0.31
ACENet[56] 338.5 0.13
MIRNet[59] 387.7 0.13
DiffLL[21] 355.1 0.36
DR2[53] 341.0 0.06
GFPGAN[50] 316.8 0.06
Ours 284.2 0.43
Table 6: Comparisons with state-of-the-art works on the real-captured test cases under extreme conditions.
  Photons Per Pixel 5 10
[…]
UTVNet[65] 11.06 0.1756 0.8275 11.71 0.2226 0.8067
ACENet[56] 9.97 0.1741 0.7949 10.81 0.2326 0.7683
SNR-aware[57] 9.11 0.0590 0.8527 9.91 0.0864 0.8366
MIRNet[59] 9.05 0.3376 0.8546 10.02 0.3896 0.8004
DiffLL[21] 11.96 0.2832 0.8448 12.67 0.3109 0.8143
DR2[53] 7.17 0.1902 0.7435 8.08 0.2399 0.7437
GFPGAN[50] 7.20 0.2211 0.7251 8.13 0.2845 0.7185
Ours 17.93 0.5375 0.4731 18.49 0.6115 0.4120
Table 7: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset under extreme conditions.
  Methods 5 PPP 10 PPP
UTVNet[65] 0.0998 0.1129
ACENet[56] 0.0999 0.1233
SNR-aware[57] 0.0988 0.1001
MIRNet[59] 0.0991 0.1747
DiffLL[21] 0.1004 0.1209
DR2[53] 0.0987 0.1043
GFPGAN[50] 0.0949 0.1176
Ours 0.5995 0.7271
Table 8: Comparisons of identity score with other methods on CelebAMask-HQ dataset under extreme conditions.
  Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
w/ Encoder 12.62 0.2574 0.6315 0.6906
w/ physical buffers 13.85 0.3753 0.6042 0.8039
Table 9: Comparison of ID physical buffers and encoder-based methods on the real-world photo test set. For a fair comparison, we use two same gallery images for each individual for both the encoder-based and physical buffer-based approaches.
Appendix D More Ablation Study
The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the User’s gallery photos. As claimed in the paper, facial features extracted from a single image are inherently limited. In Fig. 10, we evaluate the effect of the number of gallery images used to generate the ID physical buffer on identity preservation performance. We analyze it with the real-captured extreme test set where 4 gallery images were collected from the subject’s phone. The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image. Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it, which corresponds to the case with zero gallery photos.
Appendix E More Implementation Details
During training and testing, the image size is 256 $\times$ 256. The architecture of the diffusion model is adapted from [11]. We modify the architecture so that the model can take the concatenation of physical buffers and degraded images as conditions. The model parameters are configured as follows: the number of channels is set to 128, channel multiple is (1, 1, 2, 2, 4, 4), head channels are 128, attention resolution is 16, dropout rate is 0.1, diffusion steps are 1000, and the two hyperparameters P2_gamma and p2_k are both set to 1.0. For additional architectural details, refer to [11]. During inference, we adopt the DDIM method [45] with 200 denoising steps. The model is trained using the Adam optimizer [26] with a learning rate of $10^{-4}$. The weight decay is set to 0. All experiments are performed on a single NVIDIA A100 GPU.
Appendix F Encoder vs. Physical Buffer?
Many previous studies have used encoder-based methods to extract latent facial feature information from reference photos[49]. To evaluate whether the encoder can effectively extract enough facial features to guide the diffusion model in solving the low-light enhancement and denoising task, we conducted an experiment. As illustrated in Fig. 11, we utilized an encoder (ResNet-18 [18]) to extract latent codes from the target individual’s gallery photos. Following the method employed by ADM [11] for time embedding, we scaled and shifted the features in each layer of the diffusion model using the latent codes. This process integrates the identity information extracted by the encoder into the diffusion model.
We compared the restored images generated with the encoder-based method with the proposed physical buffer-guided method on our real-captured photo set. The quantitative and visual comparisons are presented in Table 9 and Fig. 15, respectively. The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the encoder-based approach. One major reason for this is that the ID physical buffers serve as robust conditions, enabling the extraction of comprehensive facial features from the gallery. In contrast, the encoder-based method is less focused on facial features and is often influenced by irrelevant factors such as background and hairstyle. Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
Appendix G Gallery Photos
Fig. 19 presents the sampled gallery photos for each test subject included in the main paper and supplementary material. Additionally, the gallery photos of the real-captured male subject are displayed in Fig. 9.
Figure 12: Visual comparison on CelebAMask-HQ test set under extreme conditions (10 PPP).
Figure 13: Visual comparison on CelebAMask-HQ test set under extreme conditions (5 PPP).
Figure 14: Visual comparison on the real-world captured SUBJECT under extreme conditions (zoomed-in area in the input is scaled for visual purposes).
Figure 15: Visual comparison on different identity representation strategies (physical buffer and face encoder).
Figure 16: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
Figure 17: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
Figure 18: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
Figure 19: Gallery samples of testing persons appeared in the main paper and supplementary materials.

### v2

@@
 Abstract
-While smartphone cameras today can produce astonishingly good photos, their performance in low light is still not completely satisfactory because of the fundamental limits in photon shot noise and sensor read noise.
-Generative image restoration methods have demonstrated promising results compared to traditional methods, but they suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low.
-Recognizing the availability of personalized photo galleries on users’ smartphones, we propose Personalized Generative Denoising (PGD) by building a diffusion model customized for different users.
-Our core innovation is an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery.
-This ID-consistent physical buffer provides a strong prior that can be integrated with the diffusion model to restore the degraded images, without the need of fine-tuning.
-Over a wide range of low-light testing scenarios, we show that PGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches.
-Figure 1: Using gallery photos from a user’s smartphone, we can restore low-light, noisy facial images.
-Our method produces finer details and better identity compared to current state-of-the-art face restoration approaches like DR2 [53] and GFP-GAN [50].
+Modern cameras produce remarkably high-quality images, yet their performance in low-light conditions remains suboptimal due to fundamental limitations in photon shot noise and sensor read noise.
+While generative image restoration methods have shown promising results compared to traditional approaches, they often suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low.
+Leveraging the availability of personalized photo galleries on users’ smartphones, we introduce Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that builds a customized diffusion model for individual users.
+Our key innovation lies in the development of an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery.
+This ID-consistent physical buffer serves as a robust prior that can be seamlessly integrated into the diffusion model to restore degraded images without the need for fine-tuning.
+Over a wide range of low-light testing scenarios, we show that DiffPGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches.
+Figure 1: Using gallery from a phone user, we can restore low-light noisy facial images.
+Our method produces finer details and better identity consistency compared to other state-of-the-art restoration approaches like MIRNet [61, 62], GFP-GAN [52], DiffLL [22] and CodeFormer [67] on both simulated (top row) and real-captured (bottom row) low-light noisy inputs.
 22footnotetext: Project page: https://genai-restore.github.io/DiffPGD/.
@@
 Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [36], total variation [44, 4], wavelets [9] to non-local means [3], BM3D [10], and recently convolution neural networks [25, 64, 50, 6, 2, 18, 57], transformers [35, 38, 63], and diffusion models [26, 60, 68].
-A common theme across existing methods is that they all focus on a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images.
+A common theme across existing methods is to find and utilize a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images.
 While they perform well in moderately difficult problems, the generality of these methods cannot be extended to heavily corrupted images.
@@
 Why Gallery Photos?
-Smartphone cameras today store hundreds if not thousands of a user’s photos, captured at different times, in different places, and under different lighting conditions.
+Smartphone cameras today typically store hundreds if not thousands of a user’s photos, captured at different times, in different places, and under different lighting conditions.
 While these images have many variations, they are all about the same person(s).
 Therefore, if the imaging goal is to take a photo of this user, the gallery on the phone would be the best source to build a prior $p(\mathbf{x})$.
-The situation is summarized in Fig.
+This concept is illustrated in Fig.
 2.
-In the context of diffusion-based image restoration, the original solution space can be large because many candidate solutions are consistent with the noisy observation.
+In the context of diffusion-based image restoration, the initial solution space can be large since many candidate solutions are consistent with the noisy observation.
 The gallery provides a strong constraint to the search problem.
-This allows us to search for better quality images with high precision of the person’s identity.
+This allows us to search for better quality images and preserve the person’s identity.
 Figure 2: The restoration of inputs degraded by noise and low-light conditions is highly ill-posed.
@@
 There are, however, two technical questions:
-  • What class of images to focus on: In this paper, we will focus on human faces because of their relevance to most smartphone users although our algorithm can be equally applied to other image classes.
+  • What class of images we should focus on: In this paper, we focus on human faces because of their relevance to phone users.
 We assume that the gallery photos have been selected and are reasonably informative.
 Product-level engineering such as pre-processing and selection of the gallery photos is important but beyond the scope of this paper.
-  • How to incorporate this class-specific $p(\mathbf{x})$ into the restoration model: The bigger (and harder) question is that given the gallery photos, how do we efficiently extract the prior information and improve the restoration?
-We assume that the gallery set is not too big and so we cannot train a restoration model from scratch.
+  • How to incorporate the class-specific $p(\mathbf{x})$ into the restoration model: The bigger (and harder) question is that given the gallery photos, how do we efficiently extract the prior information and improve the restoration?
+We do not want to train a restoration model from the scratch because the gallery can be small.
 We want to avoid fine-tuning as much as possible during inference.
-Moreover, we want to make sure that the person’s identity is preserved.
+Moreover, the person’s identity needs to be preserved.
 Physical Buffers to the Rescue.
 Given the gallery photos, what kind of prior information would be useful for restoration?
-Advancements in computer vision have made it possible to extract detailed facial physical buffers—including albedo and normal maps—from a person’s gallery of images[42, 63].
-These physical buffers capture intrinsic properties such as skin color and surface geometry, effectively encoding an individual’s unique identity and fine facial details.
-(See Fig.
-3.) At the same time, it also eliminates the influence of environmental lighting, pose, and other identity-independent variables.
-We will leverage this rich prior information to improve restoration.
-To convince readers that physical buffers could be useful to our denoising task, we show in Fig.
-3 a toy experiment.
-We consider three types of degradation: low-resolution (LR), motion blur, and noise.
-In each case, we run an existing physical buffer extraction method DECA[14] to extract the albedo and surface normal of the input.
-We observe that despite the different types of degradation, the physical buffers (albedo and surface normal) are fairly accurate.
-The only failure case is when the noise is extremely strong.
-Therefore, if we can develop solutions to manage the heavy noise case, we will be able to use physical buffers as a signal prior to guiding the denoising process.
+Advancements in computer vision have made it possible to extract facial physical buffers—including albedo and normal maps—from a person’s gallery of images[43, 65, 47].
+These physical buffers capture crucial identity-defining properties such as surface geometry and skin color, effectively encoding an individual’s unique identity.
+At the same time, it also eliminates the external influence of environmental lighting, pose, and other identity-independent variables.
+By incorporating this robust prior that disentangles the external influences, our approach can effectively extracts identity information from the gallery.
 Contributions. Our contributions are threefold:
   1. 1.
-    Personalized Low-Light Denoise and Enhancement Framework: We introduce a new framework that leverages diffusion models and person-specific priors from photo galleries to address identity loss in restored low-light facial images.
+    Personalized Low-Light Denoise and Enhancement Framework: We introduce a new framework that leverages diffusion models and person-specific priors from photo galleries to denoise low-light facial images while presering the identity.
   2. 2.
@@
   3. 3.
-    Easy Deployment: Our method operates easily without requiring fine-tuning for individual users, making it practical for real-world applications and scalable to multiple users.
+    Easy Deployment: Our method operates easily without requiring fine-tuning for individual users, making it practical and scalable to multiple users.
 Extensive experiments demonstrate that our approach outperforms existing techniques in both visual quality and identity preservation, providing a significant advancement in low-light facial image restoration.
@@
 However, to the best of our knowledge, no specific approach currently exists for joint enhancement and denoising of facial images in low-light conditions.
-Figure 3: For images affected by mild low resolution, motion blur, or noise, DECA [14] can accurately extract physical buffers.
-However, DECA fails to obtain precise physical buffers for images degraded by severe low light and noise.
+Figure 3: Path to the idea. Using facial physical buffers and gallery images can help preserve human face identity information.
 Deep Face Restoration.
@@
 In contrast, the gallery-based approach [13] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration.
+Despite the effectiveness [13] using gallery photos of users, it still requires multi-stage training and fine-tuning on each user during inference.
 3 Method
+Figure 4: For images affected by mild low-light noise, phycial buffers can be accurately extracted.
+However, when encountering the images degraded by severe low light and noise, precise physical buffers are difficult to obtained.
 3.1 Restoration by Diffusion
-A diffusion model approximates the clean image distribution $p_{\theta}(x_{0})$ by learning a model $\theta$ that reverses the noise-adding process.
-In Denoising Diffusion Probabilistic Models (DDPMs) [20], Gaussian noise is progressively added to a clean image $x_{0}$:
-$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon,\quad\text{where}\quad\epsilon\sim\mathcal{N}(0,I)$ (2)
+A diffusion model approximates the clean image distribution $p_{\theta}(\mathbf{x}_{0})$ by learning a model $\theta$ that reverses the noise-adding process.
+In Denoising Diffusion Probabilistic Models (DDPMs) [21], Gaussian noise is progressively added to a clean image $\mathbf{x}_{0}$:
+$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\text{where}\quad\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (2)
 The reverse process then removes the noise to reconstruct the clean image.
-After training, for any given time $t$ and the corresponding noisy image $x_{t}$, the model can iteratively denoise by sampling from $p(x_{0}|x_{t})$.
-In restoration tasks, like low-light denoising, the goal is to recover a high-quality image $x_{0}$ from a low-quality input $y_{0}$.
-Unlike pure generation, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $y_{0}$.
-During training, the model uses $y_{0}$ as a condition, and once the model is trained, for any given time $t$ and the corresponding noisy image $x_{t}$, it can iteratively denoise by sampling from $p(x_{0}|x_{t},y_{0})$ using the trained model.
-Figure 4: Path to the idea. Using facial physical buffers and gallery images can effectively preserve human face identity information.
+After training, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, the model can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t})$.
+In restoration tasks, like low-light denoising, the goal is to recover a high-quality image $\mathbf{x}_{0}$ from a low-quality input $\mathbf{y}$.
+Unlike pure generation, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $\mathbf{y}$.
+During training, the model uses $\mathbf{y}$ as a condition, and once the model is trained, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, it can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y})$ using the trained model.
 Figure 5: The overall architecture of the proposed method.
@@
 For a high-quality gallery, we use LAP [65] to extract the albedo and normal information for each photo and apply adaptive aggregation to fuse the entire gallery.
-The extracted albedo represents base skin color and texture details, such as skin tone, freckles, and pigmentation, while normal captures facial geometry, including wrinkles, pores, and fine surface variations.
+The extracted albedo represents base skin color and facial appearance, while normal captures facial geometry.
 In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
-3.2 Why Gallery + Physical Buffer?
+3.2 Why Combining Gallery and Physical Buffer?
 We formalize our core approach of extracting identity-consistent physical buffers from gallery photos to guide the diffusion model for restoration.
-This method is based on the observation that using facial physical buffers or gallery images can effectively preserve identity information during restoration tasks involving diffusion models.
-To evaluate the capability of physical buffers and gallery images in preserving identity information without relying on prior identity knowledge from degraded inputs, we train the following models using a blind scheme, where the degraded input ($y_{0}$) is not used as a condition during training.
-As illustrated in Fig.
-4, the models are designed as follows:
-  • Model 1: Trained exclusively on generic images as a purely generative model.
-  • Model 2: Trained on generic images while conditioned on the facial physical buffer extracted from the input image.
-This model incorporates face identity and facial features through the physical buffer.
-  • Model 3: Fine-tuned on Model 1 with the user’s gallery photos. This model learns identity features from the target person’s gallery photos during fine-tuning.
-  • Model 4: Fine-tuned on Model 2 with the user’s gallery photos.
-This model integrates facial identity and features from the physical buffer of the input image, as well as the target person’s gallery photos.
-We test these models using a simple low-light noisy input.
-Model 1 generates random faces, as expected for a purely generative model.
-In contrast, Model 2 produces images with facial features (e.g., nose, mouth, skin color) resembling those of the user.
-This is because the physical buffers extracted from the degraded user image (refer to Fig.
-3) were used as conditioning.
-These results demonstrate that facial features and identity information can be incorporated by leveraging physical buffers.
-The outputs from Model 3 are all different user faces, compared to Model 1, showing that fine-tuning with gallery photos can effectively transfer strong identity information.
-The outputs from Model 4 are the best-expected results, demonstrating the effectiveness of using gallery photos and physical buffers to maintain the person’s identity.
-However, despite the powerful combination of gallery images and physical buffer, there are limitations:
+This method is based on the observation that using facial physical buffers or gallery images can help preserve identity information during restoration tasks involving diffusion models.
+To evaluate the capability of physical buffers and gallery images, we conducted four simple experiments under a blind scheme (excluding $y_{0}$ during training), as illustrated in Fig.
+3:
+  • Exp 1: Model trained solely on generic images (purely generative).
+  • Exp 2: Model trained on generic images, conditioned on the facial physical buffer extracted from the input image.
+  • Exp 3: Model trained on generic images, then fine-tuned with the user’s gallery photos.
+  • Exp 4: Model trained on generic images, fine-tuned with the user’s gallery photos, and conditioned on the facial physical buffer.
+We tested these models on a mildly noisy input.
+Exp 1 produces random faces, as expected from a purely generative model.
+In contrast, Exp 2 uses physical buffers obtained via DECA [15], which can extract coarse face buffers from a single 2D photo (see Fig.
+4), yielding faces with user-like features (e.g., nose shape, mouth and skin tone).
+Hence, physical buffers can convey key identity cues.
+Exp 3 outputs distinct faces of the same user, compared to Exp1, showing that fine-tuning with gallery photos effectively transfers identity information.
+Finally, Exp 4 yields the best results, showcasing that the collaborative use of gallery photos and facial physical prior can achieve the optimal outcomes.
+While Fig. 3 shows us a design direction, we need to overcome:
   1. (i)
-    Fine-tuning: Extracting identity information from gallery photos via fine-tuning is time-consuming and inconvenient.
-Each time when we switch to a different user, the model must be fine-tuned again.
-Additionally, since the size of a gallery is usually much smaller than a pre-trained model, fine-tuning can easily overfit.
+    Fine-tuning: Extracting identity via fine-tuning requires significant time and must be repeated for each user.
+Additionally, the small size of a gallery compared with a large pre-trained model can lead to overfitting.
   2. (ii)
-    Extracting physical buffers from the degraded input image: If the input is heavily degraded, the physical buffers extracted from it can be wrong (refer to Fig.
-3), and result in terrible results as shown in Fig.
-4.
-Besides, the facial feature extracted from a single image is limited, not to mention from a degraded image.
-In conclusion, we want to avoid fine-tuning the model on the gallery and we do not want to extract incorrect physical buffers due to noise.
-3.3 ID-Consistent Physical Buffer
-To address these limitations, we propose extracting physical buffers directly from existing clean gallery photos and using them as conditions to guide the diffusion model during training.
-This approach offers significant benefits: clean gallery photos allow us to form comprehensive and robust identity-consistent physical buffers containing detailed facial identity features of the target person.
+    Unreliable buffers from degraded inputs: As studied in the face reconstruction literature, low-quality inputs consistently present additional challenges [15].
+Severe degradation undermines the extraction of physical buffers (Fig.
+4), resulting in poor reconstruction.
+Moreover, features captured from a single degraded photo are limited, not to mention from a degraded one.
+In conclusion, to avoid repeated fine-tuning on galleries and reduce the risk of erroneous buffers from noisy inputs, it is more desirable to jointly employing gallery photos and facial physical priors, as we will discuss below.
+3.3 Extracting Physical Buffer from Photo Gallery
+Based on the preceding analysis, we propose extracting physical buffers directly from existing clean gallery photos and using them as conditions to guide the diffusion model during training.
+This approach offers significant benefits: clean gallery photos allow us to form comprehensive and robust identity-consistent physical buffers containing more acurate facial identity features of the target person.
 Specifically, We will extract the albedo and normal maps of the target person, which are two key attributes used to describe the physical characteristics of a face.
   • Albedo captures the intrinsic color of an object’s surface—that is, the subject’s color properties without any lighting effects.
-  • Normal information is used to describe the fine bumps and details such as skin wrinkles and pores.
-By conditioning on these strong identity physical buffers, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction.
+It reveals a person’s facial appearance, e.g.
+eyebrows, lip coloration, and other distinct texture elements can reflect individual identity traits.
+  • Normal encodes the surface orientation at each pixel, which provides hints about the underlying shape characteristics.
+By conditioning on these identity physical buffers, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction.
 Consequently, the model is better equipped to preserve the target person’s identity and facial features, even under severe degradation.
-3.4 Extracting Physical Buffers
-Extracting physical buffers for a person from a set of gallery photos is a more challenging task than extracting these buffers from a single input, as in methods like DECA.
+However, extracting physical buffers for a person from a set of gallery photos is a more challenging task than extracting these buffers from a single input, as in methods like DECA.
 This complexity arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses.
-To ensure these buffers comprehensively represent the person’s identity and facial details, they must be aligned to maintain identity consistency.
-Fortunately, existing works have addressed this challenge.
+To ensure these buffers comprehensively represent the person’s identity and facial properties, they must be aligned to maintain identity consistency.
 As shown in the physical buffer extractor module in Fig.
-5, we adopt the aggregation net design from the LAP [64] to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
+5(a), we adopt the aggregation net design from the LAP [65] to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
 3.4 Model Framework
@@
 5, the model leverages ID physical buffers (albedo $a^{i}$ and normal $n^{i}$) extracted from person i’s gallery photos as conditions to guide our diffusion model during training directs the generative process toward a personalized space.
-Within the aggregation network, it contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting consistent face.
-To model albedo and normal, two separate aggregation networks denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed.
-Given a photo gallery of N images $\{\textbf{I}_{k}^{i}\}_{k=1}^{N}$ of person i, each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $x_{k}^{a}$ and $x_{k}^{n}$, respectively.
-To derive a global representation of the identity for person i based on $\{x_{k}^{a},x_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted, This method learns a weight set $\{w_{k}^{a},w_{k}^{n}\}_{k=1}^{N}$ according to the quality of each image in$\{\textbf{I}_{k}^{i}\}_{k=1}^{N}$.
-The combined global ID-code $x_{c}^{a}$ and $x_{c}^{n}$ for texture and depth are calculated as:
-$x_{c}^{a}=\sum_{k=1}^{N}w_{k}^{a}x_{k}^{a},\quad x_{c}^{n}=\sum_{k=1}^{N}w_{k}^{n}x_{k}^{n}.$ (3)
-Then the global ID-code $x_{c}^{a}$ and $x_{c}^{n}$ are then passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $a^{i}$ and normal $n^{i}$.
-Additionally, with the low-light noisy image $y_{0}^{i}$ as another condition during training, other features, such as pose, hairstyle, and background inside the image beyond facial features can be effectively captured as well.
-After the model is trained, during testing, the process begins with a Gaussian noise image $x_{T}$, which is iteratively transformed into the clean, enhanced image $x_{0}^{i}$ by sampling from $p(x_{0}^{i}|x_{T},y_{0}^{i},a_{i},n_{i})$ using the trained model.
+The aggregation network contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting the consistent representation.
+To model albedo and normal, two separate aggregation networks, denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed.
+Given a photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i, each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $\mathbf{\alpha}_{k}^{a}$ and $\mathbf{\alpha}_{k}^{n}$, respectively.
+To derive a global representation of the identity for person i based on $\{\mathbf{\alpha}_{k}^{a},\mathbf{\alpha}_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted, This method learns a weight set $\{\mathbf{w}_{k}^{a},\mathbf{w}_{k}^{n}\}_{k=1}^{N}$ according to the quality of each image in$\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$.
+The combined global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ for texture and depth are calculated as:
+$\mathbf{\alpha}_{c}^{a}=\sum_{k=1}^{N}\mathbf{w}_{k}^{a}\mathbf{\alpha}_{k}^{a},\text{ and}\quad\mathbf{\alpha}_{c}^{n}=\sum_{k=1}^{N}\mathbf{w}_{k}^{n}\mathbf{\alpha}_{k}^{n}.$ (3)
+Once calculated, the global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ are passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$.
+During inference, the process begins with a Gaussian noise image $\mathbf{x}_{T}^{i}$, which is iteratively transformed into the clean, enhanced image $\mathbf{x}_{0}^{i}$ by sampling from $p(\mathbf{x}_{0}^{i}|\mathbf{x}_{T}^{i},\mathbf{y}^{i},\mathbf{a}^{i},\mathbf{n}^{i})$ using the trained model.
+With the low-light noisy image $\mathbf{y}^{i}$ as another condition during training, other features, such as pose, hairstyle, and background inside the image beyond facial features can be effectively captured as well.
 Our experiments demonstrate that identity information is effectively preserved with the guidance of these ID-consistent buffers, enabling our model to achieve superior results in human face restoration.
 4 Experiments
+Figure 6: Visual comparison on real cases.
+The first column presents the input images along with their capture parameters.
+In the second column, we enhance the brightness to better reveal the noise.
+Please zoom in for the best visual experience.
 4.1 Dataset
@@
 Refer to supplementary for more details on the simulation.
-Testing Data.
-For testing DiffPGD, we utilize both simulated and real-world low-light noisy data.
+  Photons Per Pixel 5 10 13 26
+Methods ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$
+SNR-aware[59] 0.2156 188.3311 0.1302 8.4541 0.4240 139.8193 0.0786 9.5462 0.4999 124.1162 0.0626 10.1712 0.6937 99.9875 0.0383 12.4961
+MIRNet[61, 62] 0.1229 294.7381 0.2414 8.6008 0.2860 214.7785 0.1534 9.8572 0.3830 187.2571 0.1286 10.4209 0.6212 136.1320 0.0793 12.8992
+DiffLL[22] 0.2910 167.3905 0.1068 17.3506 0.4544 150.4559 0.1117 18.3360 0.5095 140.2738 0.0996 19.9547 0.6833 119.2693 0.0760 23.3180
+GFPGAN[52] 0.1941 388.9293 0.4256 7.1990 0.2075 233.3173 0.1521 8.1283 0.2976 190.7264 0.1326 9.5483 0.5976 158.9475 0.0829 11.4855
+CodeFormer[67] 0.2863 111.4544 0.04908 18.2406 0.4940 91.1244 0.0326 19.2692 0.5563 84.7955 0.0281 20.2306 0.7338 80.6582 0.0263 24.5512
+Ours 0.3714 77.5743 0.0084 17.9334 0.5225 71.8645 0.0063 18.4999 0.5938 69.5567 0.0042 19.0910 0.7118 59.1975 0.0030 23.7226
+Table 1: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset on 5PPP, 10PPP, 13PPP, and 26PPP. The lower the PPP, severe the degradation is.
+Testing Data. For testing DiffPGD, we utilize both simulated and real-world captured low-light noisy data.
 The simulation setup assumes a $12$-bit image sensor with parameters similar to that of Canon DSLR cameras: a Full Well Capacity of $36000\text{e}^{-}$, Quantum Efficiency of $0.42$, dark current of $7\text{e}^{-}$, and read noise of $6\text{e}^{-}$.
 These parameters ensure realistic noise characteristics for evaluating model performance.
-We collect the real-world test data using the Sony $\alpha 6400$ DSLR camera at an approximate light intensity of $0.13$ Lux with an ISO of $25600$ and an exposure time of $\frac{1}{8}$ seconds.
+We also captured 100 RAW-format face images as real-world test dataset using a Sony $\alpha 6400$ DSLR camera paired with a Sony 18-135mm lens at night.
+These images feature diverse low-light and dark backgrounds, various angles and poses, as well as different focal lengths, apertures, ISO values, and shutter speeds to reflect the rich characteristics of real-world low-light photography.
+We cropped these images into 256×256 face patches.
 4.2 Implementation Details
@@
 Within each ID, there are 3–20 images.
+The number of gallery images we used per person varies from 3 to 6, with most around 5.
 The Adam optimizer is applied with a learning rate of 1e-4.
 Training is performed for 200,000 iterations with a batch size of 32, resulting in a total of 6,400,000 samples seen by the model.
-Figure 6: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
-Figure 7: Visual comparison on real-captured photo test (zoomed-in area in the input is scaled for visual purposes).
-Our method outputs the best denoised and enhanced result, and successfully restores fine facial details, such as moles, wrinkles.
+4.3 Performance in Real-Captured Scenarios
+Testing model performance in real-life low-light scenarios is crucial and challenging, as real-world conditions often involve uncontrollable factors compared to simulated ones.
+Since no existing data set addresses the restoration of real-life low-light noisy face photos, we captured photos of real persons under indoor and outdoor low-light scenarios , as illustrated in Section 4.1.
+We compare our methods with several state-of-the-art works: UTV-Net[66], GFPGAN[52], SNR-aware[59], MIRNet[61, 62], DiffLL[22], Codeformer [67].
+These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research.
+Particularly, GFPGAN and Codeformer are designed specifically for face image restoration, which is highly competitive with our method.
+We re-trained/fine-tuned all the baselines with the same training dataset as ours for fair comparison.
+Quantitative and qualitative results are presented.
+To assess the quality of restored images, we employ the Fréchet Inception Distance (FID) [20] and the Kernel Inception Distance (KID) [1], as these metrics offer a robust evaluation aligned with human perception.
+Since ground-truth images are unavailable for real-captured cases, we compute the FID and KID score between the model’s output and the corresponding subjects’ gallery photos.
+As shown in Fig.
+6, our method produces images with the highest quality and level of detail.
+The restored images exhibit more natural skin tones and well-preserved facial identity features, such as eye shape, eyebrows, and lips, while maintaining the highest consistency in facial identity.
+Since our work focuses on face images, we also check on the identity preservation performance.
+Besides the image quality metrics, we adopt the identity score which calculates the cosine similarity of the image features extracted by ArcFace[11] to further evaluate the identity consistency.
+Again, since ground truth images are unavailable for real-captured cases, we calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo of the corresponding subjects.
+The results are also shown in Table 2.
+Again, our method has the highest identity score, and the visual results are consistent with the quantitative metric results.
+The results demonstrate the supreme ability of our proposed DiffPGD model on real low light denoising and enhancement task.
 4.4 Performance in Simulated Scenarios
-In simulated scenarios, we compare our methods with several state-of-the-art works: UTVNet[65], ACENet[56], GFPGAN[50], SNR-aware[57], DR2[53], MIRNet[59] and DiffLL[21].
-These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research.
-Particularly, GFPGAN and DR2 are designed specifically for face image restoration, which is highly competitive with our method.
-We use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set.
-We simulated five low-light noisy levels by setting different photons per pixel (PPP), which range from 13 PPP to 65 PPP, the lower the PPP, severe the degradation is.
-Quantitative and qualitative results are provided.
-Firstly, we use PSNR, SSIM, and the perceptual metric (LPIPS[62]) to measure the restored image quality.
-The results are shown in Table 1, it can be seen that our method has surpassed all previous methods by a large scale.
-Demonstrating the supreme ability of our proposed DiffPGD model on low light denoising and enhancement problems.
-Since our work focuses on face images, we also check on the identity preservation performance.
-Therefore, besides the PSNR, SSIM, and LPIPS metrics, we adopt the identity score which calculates the cosine similarity of the image features extracted by ArcFace[10] to further evaluate the identity consistency.
-The results are shown in Table 2.
-Again, our method has the highest identity score.
-The advantage of our model on preserving the identity is especially outstanding in heavily degraded cases.
-Visualized comparisons are presented in Fig.
-6.
+In simulated scenarios, We use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set.
+We simulated various low-light noisy levels by setting different photons per pixel (PPP), i.e., the average number of photons arriving at the sensor within the exposure time.
+The lower the PPP, severe the degradation is.
+A lower photon level increases the impact of read noise and dark current on the final output, thereby decreasing the quality of the reconstructed images.
+Quantitative results are provided in Table 1.
 The image restored from our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistent best.
-Again, the visual results are consistent with the quantitative metric results, the advantage of our model on preserving the identity is especially outstanding in the heavily degraded cases.
-The sampled gallery photos of the test subject can be found in Fig.
-19.
-Refer to Sec.
-B in the supplementary material for more visual results.
-  Photons Per Pixel 13 26 39 52 65
-[…]
-UTVNet[65] 12.08 0.2644 0.7816 13.26 0.3685 0.7169 13.99 0.4332 0.6754 14.50 0.4783 0.6419 14.87 0.5118 0.6148
-ACENet[56] 11.25 0.2815 0.7431 12.27 0.4022 0.6725 13.19 0.4670 0.6276 13.87 0.5120 0.5894 14.43 0.5479 0.5577
-SNR-aware[57] 10.74 0.1186 0.8193 11.99 0.2031 0.7650 12.34 0.2546 0.7230 12.53 0.2935 0.6852 12.76 0.3271 0.6507
-MIRNet[59] 11.64 0.4553 0.7476 13.64 0.5416 0.6491 13.90 0.5724 0.5874 14.12 0.5960 0.5355 14.323 0.6154 0.4934
-DiffLL[21] 13.30 0.3472 0.7872 14.66 0.4282 0.7189 15.47 0.4588 0.6736 15.96 0.4805 0.6387 16.25 0.4969 0.6072
-DR2[53] 13.50 0.3448 0.7573 15.16 0.5141 0.6430 16.22 0.5655 0.5871 16.85 0.5909 0.5503 17.54 0.6209 0.5181
-GFPGAN[50] 13.53 0.4256 0.7049 15.56 0.5446 0.5713 16.60 0.5960 0.5156 17.41 0.6322 0.4748 17.79 0.6563 0.4431
-Ours 21.10 0.6287 0.3862 22.80 0.6914 0.3394 23.85 0.7195 0.3156 24.11 0.7354 0.2965 24.38 0.7399 0.2947
-Table 1: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset.
-  Methods 13 PPP 26 PPP 39 PPP 52 PPP 65 PPP
-UTVNet[65] 0.1572 0.3654 0.5201 0.6197 0.6938
-ACENet[56] 0.1734 0.4173 0.5769 0.6756 0.7240
-SNR-aware[57] 0.1069 0.1921 0.3114 0.4171 0.4971
-MIRNet[59] 0.3071 0.5613 0.6694 0.7246 0.7794
-DiffLL[21] 0.1751 0.3779 0.5033 0.6107 0.6682
-DR2[53] 0.2456 0.5128 0.6358 0.6905 0.7239
-GFPGAN[50] 0.2506 0.5392 0.6569 0.7212 0.7615
-Ours 0.5995 0.7271 0.7702 0.8044 0.8253
-Table 2: Comparisons of identity score with other methods on CelebAMask-HQ dataset.
-  Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
-UTVNet[65] 12.02 0.3901 0.6954 0.8062
-ACENet[56] 11.36 0.3962 0.6817 0.8125
-SNR-aware[57] 10.18 0.1922 0.7274 0.6111
-MIRNet[59] 12.48 0.4315 0.7132 0.6914
-DiffLL[21] 13.18 0.3899 0.7551 0.7686
-DR2[53] 14.08 0.1829 0.7350 0.5309
-GFPGAN[50] 14.18 0.3714 0.6784 0.7138
-Ours 14.02 0.4362 0.5954 0.8248
-Table 3: Comparisons with state-of-the-art works on real-captured photo test.
-  Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
-w/o physical buffers 12.96 0.3801 0.6196 0.7538
-w/ physical buffers 14.02 0.4362 0.5954 0.8248
+(Refer to supplementary for more visual comparisons).
+From the table, we could see that CodeFormer surpasses us on PSNR.
+Fig.
+7 illustrates a visual comparison between our method and CodeFormer under inputs ranging from 5 to 65 PPP.
+We observe that CodeFormer tends to generate smoother images while we could preserve better high-frenquency details.
+Furthermore, our method demonstrates superior identity preservation even at low resolutions, such as 5 PPP to 13 PPP, while exhibiting fewer artifacts.
+The advantage of our model in maintaining identity consistency is particularly evident in heavily degraded cases.
+The results highlight the strong performance of our proposed DiffPGD model in low-light denoising and enhancement tasks.
+  Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
+SNR-aware[55] 252.2397 0.1264 0.1990
+UTVNet[66] 341.8502 0.1617 0.2002
+MIRNet[61, 62] 276.3853 0.1659 0.2210
+DiffLL[22] 252.1553 0.1421 0.3196
+GFPGAN[52] 252.7411 0.1249 0.1525
+CodeFormer[67] 281.9744 0.1459 0.2965
+Ours 251.9318 0.0882 0.3708
+Table 2: Comparisons with state-of-the-art works on the real-captured test cases.
+  Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
+w/o physical buffers 272.6148 0.1281 0.3169
+w/ physical buffers 251.9318 0.0882 0.3708
 Table 3: Ablation comparisons of using ID physical buffers on real-captured photo test.
-4.4 Performance in Real-Life Scenarios
-Testing model performance in real-life low-light scenarios is crucial, as real-world conditions often involve uncontrollable factors compared to simulated ones.
-Since no existing dataset addresses the restoration of real-life low-light noisy face photos, we created a custom test set.
-Specifically, we printed five photos of five target individuals and captured images of these photos under low-light conditions using a camera with the settings specified in Section 4.1.
-As shown in Fig.
-7, our model demonstrates significantly better recovery of real-captured low-light noisy inputs compared to existing models.
-In Table 3, we quantitatively compare the performance of our model with other baselines on this real-captured photo test.
-We observe that all our metrics are better except for GFPGAN’s PSNR.
-However, when looking at the visual comparison in Fig.
-7, we realize that GFPGAN recovers smoother features but loses the true facial details, resulting in a relatively fake-looking face.
-For further verification, we also photographed real individuals under low-light conditions with the camera.
-As shown in Fig.
-8, DiffPGD outperforms existing methods by a large margin.
-These results confirm that DiffPGD effectively leverages prior information from a target person’s gallery photos to guide the diffusion model in learning a personalized generative space.
-During restoration, the model accurately directs the output toward this clean personalized space.
-The gallery photos of the real-captured male subject are displayed in Fig.
-9.
-Other test subjects’ gallery photos can be found in Fig.
-19.
-Figure 8: (Left) Environmental setup, real-life low-light noisy photo taken in the dark room. (Right) Visual comparison on the real-world captured SUBJECT.
-Figure 9: Male SUBJECT’s gallery (face areas are cropped from the subject’s photos).
+Figure 7: Visual comparison on the simulated cases. Our method maintains good identity consistency even at extremely low PPP and exhibits fewer artifacts.
+Figure 8: The relationship between the identity (ID) score and FID of the restored images and the number of gallery images used to generate the ID physical buffer.
 5 Ablation Study
 ID-Consistent Physical Buffers for Identity Preservation.
-We have shown that the proposed model DiffPGD exhibits outstanding performance in both simulated and real-world testing scenarios.
-In Table 4, we compare the quantitative results of training our model with and without ID physical buffers in real-world test cases.
+We have shown that the proposed model DiffPGD exhibits outstanding performance in both simulated and real-captured testing scenarios.
+In Table 3, we compare the quantitative results of training our model with and without ID physical buffers in real-captured test cases.
 The results demonstrate that using physical buffers largely increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
+Effect of the Number of Gallery Images.
+The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the User’s gallery photos.
+As claimed in the paper, facial features extracted from a single image are inherently limited.
+In Fig.
+8, we evaluate the effect of the number of gallery images used to generate the ID physical buffer.
+We analyze it with the real-captured set where 6 gallery images were collected from each subject.
+The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image.
+Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it, which corresponds to the case without gallery photos.
 6 Conclusion
-In this work, we addressed the challenges of low-light image restoration by introducing Personalized Generative Denoising (PGD), a novel approach that leverages user-specific photo galleries to improve denoising performance.
+In this work, we addressed the challenges of low-light facial image restoration by introducing Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that leverages user-specific photo galleries to improve denoising performance.
 By incorporating an identity-consistent physical buffer, we provided a strong prior that enhances the diffusion model’s ability to restore degraded images while avoiding the need for model fine-tuning.
-Our experiments demonstrated that PGD outperforms existing diffusion-based denoising methods, delivering superior results across various testing scenarios.
-This approach not only mitigates false content generation in low-SNR conditions but also bridges the gap in low-light imaging quality, paving the way for personalized and more robust image restoration techniques.
-Acknowledgment This research was supported, in part, by a research contract from Samsung Research America, the National Science Foundation (NSF), and in part by the Intelligence Advanced Research Projects Activity (IARPA) under Contract No.
-2022-21102100004, and in part by the National Science Foundation under the grants CCSS-2030570 and IIS-2133032.
-The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of IARPA, or the U.S.
-Government.
-The U.S.
-Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
+Our experiments, which included extensive evaluations on multiple challenging datasets, demonstrated that DiffPGD not only outperforms existing diffusion-based denoising methods but also delivers superior quantitative and qualitative results across various testing scenarios.
+This approach effectively mitigates false content generation in low-SNR conditions while bridging the gap in low-light imaging quality, ultimately paving the way for personalized and more robust image restoration techniques.
 Supplementary Material
@@
 In Fig.
-16, Fig.
-17 and Fig.18, we present additional visual comparisons between our model and baseline methods.
+12, Fig.
+14, and Fig.15, we present additional visual comparisons between our model and baseline methods on real-captured and simulated cases.
 Our results consistently demonstrate superior performance, particularly in cases of severe degradation.
-Appendix C Results in More Severe Cases
-To further evaluate the ability of our model to denoise and enhance extremely degraded photos, we tested it under highly challenging conditions, including both simulated and real-life scenarios.
-For simulated cases, We simulate severely degraded test cases with 10 PPP and 5 PPP.
-As shown in Fig.12 and Fig.13, our model can still perform well in terms of enhancing and denoising the degraded inputs, whereas other models struggled significantly.
-Besides, due to the extremely low signal level in the input image, we can see some hallucination effect in the hairstyle of the generated outputs.
-Despite this, thanks to the ID physical buffer extracted from the gallery of the target person, our model could still effectively restore the facial structure of the target person.
-Quantitative comparisons are presented in Tables 7 and 8.
-We also evaluated the model on real-captured photos.
-Fig.
-14 is a visual example of the male subject taken under extremely low light conditions with high noise levels.
-From the figure, we can see that our proposed model outperforms all baseline methods in both denoising and enhancement.
-Additionally, the blind-based restoration methods DR2 and GFPGAN completely failed to produce meaningful results due to the severely low signal levels.
-This supports our observation that blind restoration approaches struggle with images degraded by extremely low light and noise.
-For these real-captured extreme test cases, we further conducted quantitative evaluations of both image quality and identity preservation, as summarized in Table 6.
-Since ground truth images are unavailable for real-captured cases, we computed the Fréchet Inception Distance(FID)[19] between the model’s output images and the subject’s gallery photos, which are 4 clean natural photos from the subject’s phone (shown in Fig.9).
-We calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo.
-The results indicate that our model consistently achieves superior image quality and identity preservation compared to other methods.
-Figure 10: The relationship between the identity (ID) score of the restored images and the number of gallery images used to generate the ID physical buffer.
 Figure 10: A network architecture diagram for identity feature restoration using a face encoder.
-  Methods FID $\downarrow$ ID Score $\uparrow$
-UTVNet[65] 293.2 0.31
-ACENet[56] 338.5 0.13
-MIRNet[59] 387.7 0.13
-DiffLL[21] 355.1 0.36
-DR2[53] 341.0 0.06
-GFPGAN[50] 316.8 0.06
-Ours 284.2 0.43
-Table 6: Comparisons with state-of-the-art works on the real-captured test cases under extreme conditions.
-  Photons Per Pixel 5 10
-[…]
-UTVNet[65] 11.06 0.1756 0.8275 11.71 0.2226 0.8067
-ACENet[56] 9.97 0.1741 0.7949 10.81 0.2326 0.7683
-SNR-aware[57] 9.11 0.0590 0.8527 9.91 0.0864 0.8366
-MIRNet[59] 9.05 0.3376 0.8546 10.02 0.3896 0.8004
-DiffLL[21] 11.96 0.2832 0.8448 12.67 0.3109 0.8143
-DR2[53] 7.17 0.1902 0.7435 8.08 0.2399 0.7437
-GFPGAN[50] 7.20 0.2211 0.7251 8.13 0.2845 0.7185
-Ours 17.93 0.5375 0.4731 18.49 0.6115 0.4120
-Table 7: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset under extreme conditions.
-  Methods 5 PPP 10 PPP
-UTVNet[65] 0.0998 0.1129
-ACENet[56] 0.0999 0.1233
-SNR-aware[57] 0.0988 0.1001
-MIRNet[59] 0.0991 0.1747
-DiffLL[21] 0.1004 0.1209
-DR2[53] 0.0987 0.1043
-GFPGAN[50] 0.0949 0.1176
-Ours 0.5995 0.7271
-Table 8: Comparisons of identity score with other methods on CelebAMask-HQ dataset under extreme conditions.
   Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
@@
 w/ physical buffers 13.85 0.3753 0.6042 0.8039
-Table 9: Comparison of ID physical buffers and encoder-based methods on the real-world photo test set.
+Table 6: Comparison of ID physical buffers and encoder-based methods on the real-captured photo test set.
 For a fair comparison, we use two same gallery images for each individual for both the encoder-based and physical buffer-based approaches.
-Appendix D More Ablation Study
-The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the User’s gallery photos.
-As claimed in the paper, facial features extracted from a single image are inherently limited.
-In Fig.
-10, we evaluate the effect of the number of gallery images used to generate the ID physical buffer on identity preservation performance.
-We analyze it with the real-captured extreme test set where 4 gallery images were collected from the subject’s phone.
-The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image.
-Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it, which corresponds to the case with zero gallery photos.
-Appendix E More Implementation Details
+Appendix C More Implementation Details
 During training and testing, the image size is 256 $\times$ 256.
@@
 For additional architectural details, refer to [12].
-During inference, we adopt the DDIM method [45] with 200 denoising steps.
+During inference, we adopt the DDIM method [46] with 100 denoising steps.
 The model is trained using the Adam optimizer [27] with a learning rate of $10^{-4}$.
@@
 All experiments are performed on a single NVIDIA A100 GPU.
-Appendix F Encoder vs. Physical Buffer?
+Appendix D Encoder vs. Physical Buffer?
 Many previous studies have used encoder-based methods to extract latent facial feature information from reference photos[51].
@@
 As illustrated in Fig.
-11, we utilized an encoder (ResNet-18 [18]) to extract latent codes from the target individual’s gallery photos.
+10, we utilized an encoder (ResNet-18 [19]) to extract latent codes from the target individual’s gallery photos.
 Following the method employed by ADM [12] for time embedding, we scaled and shifted the features in each layer of the diffusion model using the latent codes.
@@
 We compared the restored images generated with the encoder-based method with the proposed physical buffer-guided method on our real-captured photo set.
+This is a special semi-real set, in which we take photos of some printed paper photos at night.
+We prepare this semi-real set because The ground-truth will be available to be compared with.
 The quantitative and visual comparisons are presented in Table 6 and Fig.
-15, respectively.
+11, respectively.
 The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the encoder-based approach.
@@
 Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
-Appendix G Gallery Photos
-Fig.
-19 presents the sampled gallery photos for each test subject included in the main paper and supplementary material.
-Additionally, the gallery photos of the real-captured male subject are displayed in Fig.
-9.
-Figure 12: Visual comparison on CelebAMask-HQ test set under extreme conditions (10 PPP).
-Figure 13: Visual comparison on CelebAMask-HQ test set under extreme conditions (5 PPP).
-Figure 14: Visual comparison on the real-world captured SUBJECT under extreme conditions (zoomed-in area in the input is scaled for visual purposes).
+Appendix E Gallery Photos
+Fig. 13 presents the the gallery photos for the test subjects included in the Fig. 6 of the main paper.
 Figure 11: Visual comparison on different identity representation strategies (physical buffer and face encoder).
-Figure 16: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
-Figure 17: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
-Figure 18: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
-Figure 19: Gallery samples of testing persons appeared in the main paper and supplementary materials.
+Figure 12: Additional real-captured visual comparison.
+Figure 13: Gallery photos.
+Figure 14: Additional visual comparison.
+Figure 15: Additional visual comparison.

### v3

@@
 Abstract
-Modern cameras produce remarkably high-quality images, yet their performance in low-light conditions remains suboptimal due to fundamental limitations in photon shot noise and sensor read noise.
-While generative image restoration methods have shown promising results compared to traditional approaches, they often suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low.
-Leveraging the availability of personalized photo galleries on users’ smartphones, we introduce Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that builds a customized diffusion model for individual users.
+Modern cameras’ performance in low-light conditions remains suboptimal due to fundamental limitations in photon shot noise and sensor read noise.
+Generative image restoration methods have shown promising results compared to traditional approaches, but they suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low.
+Leveraging the availability of personalized photo galleries of the users, we introduce Diffusion-based Personalized Generative Denoising (DiffPGD), a new approach that builds a customized diffusion model for individual users.
 Our key innovation lies in the development of an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery.
@@
 Over a wide range of low-light testing scenarios, we show that DiffPGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches.
-Figure 1: Using gallery from a phone user, we can restore low-light noisy facial images.
-Our method produces finer details and better identity consistency compared to other state-of-the-art restoration approaches like MIRNet [61, 62], GFP-GAN [52], DiffLL [22] and CodeFormer [67] on both simulated (top row) and real-captured (bottom row) low-light noisy inputs.
-22footnotetext: Project page: https://genai-restore.github.io/DiffPGD/.
-1 Introduction
+Our project page can be found at https://genai-restore.github.io/DiffPGD/.
+I Introduction
 With the astonishing development of smartphones, it is safe to say that smartphone cameras today have surpassed digital single-lens reflex (DSLR) cameras in both popularity and functionality.
 However, the small form factor of smartphone cameras puts a tight constraint on the aperture size, hence limiting the amount of light a CMOS pixel can detect.
-In a low-light imaging environment, this creates a fundamental limitation to how short the minimum exposure needs to be and how much signal-to-noise ratio (SNR) the sensor can support.
-While some of the noise seen at low light can be mitigated using better CMOS technology (e.g., correlated double sampling to reduce the read noise [41] and deep-well pump gate to reduce the stray capacitance between the floating diffusion and the transfer gate [39, 40]), the Poisson / Bose-Einstein statistics due to random photon arrivals is a problem created by mother nature that cannot be solved by even the ideal sensors.
+In a low-light imaging environment, this creates a fundamental limitation to how short the exposure can be and how much signal-to-noise ratio (SNR) the sensor can support.
+While some of the noise seen at low light can be mitigated using better CMOS technology (e.g., correlated double sampling to reduce the read noise [Nakamura2021sensor] and deep-well pump gate to reduce the stray capacitance [ma2015pump, Ma2022qis]), the Poisson statistics due to random photon arrivals is a problem created by mother nature that cannot be solved by even the ideal sensors.
 Therefore, image denoising and enhancement by exploiting the internal structures of images become a necessary task for all smartphone cameras.
 From a signal-processing perspective, degradation caused by low light can be roughly approximated via
-$\mathbf{y}=\text{ADC}\big{(}\text{Poisson}(\alpha\mathbf{x}+\eta)\big{)}+\epsilon,$ (1)
-where $\mathbf{x}$ is the unknown clean image, $\alpha$ is the sensor gain, $\eta$ is the dark current, and $\epsilon$ is the read noise.
+$\mathbf{y}=\text{ADC}\big(\text{Poisson}(\alpha\mathbf{x}+\eta)\big)+\text{Gauss}(0,\sigma_{\text{read}}^{2}),$ (1)
+where $\mathbf{x}$ is the unknown clean image, $\alpha$ is the quantum efficiency, $\eta$ is the dark current, $\mathrm{ADC}(\cdot)$ models the analog-to-digital conversion that quantizes the sensor’s analog voltage into digital pixel values, and $\text{Gauss}(0,\sigma_{\text{read}}^{2})$ is the read noise.
 The inverse problem associated with low-light denoising is to recover $\mathbf{x}$ from $\mathbf{y}$.
-Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [36], total variation [44, 4], wavelets [9] to non-local means [3], BM3D [10], and recently convolution neural networks [25, 64, 50, 6, 2, 18, 57], transformers [35, 38, 63], and diffusion models [26, 60, 68].
+Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [Lim1989TwoDimensionalSA], total variation [Rudin1994TV, Chambolle2004tv], wavelets [Coifman1995] to non-local means [Buades2005nl], BM3D [Dabov2007BM3D], and recently convolution neural networks [Jin2017cnn, Guo2019real, wei2020physics, chi2020dynamic, Elgendy2021LowLightDA], transformers [Liang2021swinir, Lu2022low, zhang2024xformer], and diffusion models [kawar2022denoising, zhu2023DiffPIR, Wang2025GenerativePB, Yuan2025iHDRIH].
 A common theme across existing methods is to find and utilize a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images.
 While they perform well in moderately difficult problems, the generality of these methods cannot be extended to heavily corrupted images.
-Particularly in the context of human face recovery, the restored results often lack real identity and exhibit artifacts.
+In the context of human face recovery, the restored results often lack real identity and exhibit artifacts.
 This is attributed to the ill-posed nature of the problem and the lack of appropriate constraints.
@@
 This concept is illustrated in Fig.
-2.
-In the context of diffusion-based image restoration, the initial solution space can be large since many candidate solutions are consistent with the noisy observation.
-The gallery provides a strong constraint to the search problem.
+1.
+In this paper, we are interested in diffusion-based image restoration.
+The role of gallery is to provide a strong constraint to the search problem.
 This allows us to search for better quality images and preserve the person’s identity.
@@
 Product-level engineering such as pre-processing and selection of the gallery photos is important but beyond the scope of this paper.
-  • How to incorporate the class-specific $p(\mathbf{x})$ into the restoration model: The bigger (and harder) question is that given the gallery photos, how do we efficiently extract the prior information and improve the restoration?
-We do not want to train a restoration model from the scratch because the gallery can be small.
-We want to avoid fine-tuning as much as possible during inference.
+  • How to incorporate the class-specific $p(\mathbf{x})$ into the restoration model: This is a bigger (and harder) question.
+Given the gallery photos, how do we efficiently extract the prior information and improve the restoration?
+We do not want to train a restoration model from scratch because the gallery can be small.
+We also want to avoid fine-tuning during inference.
 Moreover, the person’s identity needs to be preserved.
@@
 Given the gallery photos, what kind of prior information would be useful for restoration?
-Advancements in computer vision have made it possible to extract facial physical buffers—including albedo and normal maps—from a person’s gallery of images[43, 65, 47].
+Advancements in computer vision have made it possible to extract facial physical buffers—including albedo and normal maps—from a person’s gallery of images[roth2016adaptive, Zhang2021lap, suwajanakorn2015makes].
 These physical buffers capture crucial identity-defining properties such as surface geometry and skin color, effectively encoding an individual’s unique identity.
 At the same time, it also eliminates the external influence of environmental lighting, pose, and other identity-independent variables.
-By incorporating this robust prior that disentangles the external influences, our approach can effectively extracts identity information from the gallery.
-Contributions. Our contributions are threefold:
+By incorporating this robust prior that disentangles the external influences, our approach can effectively extract identity information from the gallery.
+Contributions. Our contributions are as follows:
   1. 1.
-    Personalized Low-Light Denoise and Enhancement Framework: We introduce a new framework that leverages diffusion models and person-specific priors from photo galleries to denoise low-light facial images while presering the identity.
+    We introduce a personalized low-light denoising framework based on diffusion models.
+The proposed framework extracts person-specific priors from gallery images and integrates them into the diffusion process to effectively restore low-light facial images.
+Our solution offers three advantages: (i) We outperform existing methods in low-light conditions.
+(ii) We preserve person identity better than existing methods.
+(iii) We do not need to fine-tune for individuals.
   2. 2.
-    Physical Identity Buffers: We propose the use of physical identity buffers from existing face galleries to enrich the restoration process, enabling accurate reconstruction of human faces under severe degradation.
-  3. 3.
-    Easy Deployment: Our method operates easily without requiring fine-tuning for individual users, making it practical and scalable to multiple users.
-Extensive experiments demonstrate that our approach outperforms existing techniques in both visual quality and identity preservation, providing a significant advancement in low-light facial image restoration.
-2 Related Works
+    We propose to leverage physical buffer as an explicit prior for identity-preserving restoration.
+By learning the physical prior from clean data rather than noisy observations, the model captures intrinsic and noise-independent identity features.
+This physically grounded prior demonstrates strong generalization capability—models trained on synthetic data can effectively transfer to real-world scenarios.
+II Related Works
 Low-light Image Enhancement and Denoising.
 Low-light images often suffer from significant noise due to low signal-to-noise ratios (SNR), in addition to reduced brightness and contrast.
-Traditional low-light denoising techniques often focus on exploiting image structures, ranging from early methods like Wiener filtering [36], total variation [44, 4], and wavelets [9] to non-local means [3] and BM3D [10].
-More recently, deep learning approaches, including convolutional neural networks [25, 64, 50, 6, 2, 18, 57, 66], transformers [35, 38, 63], and diffusion-based models [26, 60, 68, 42], have emerged as promising solutions for low-light denoising.
+Traditional low-light denoising techniques often focus on exploiting image structures, ranging from early methods like Wiener filtering [Lim1989TwoDimensionalSA], total variation [Rudin1994TV, Chambolle2004tv], and wavelets [Coifman1995] to non-local means [Buades2005nl] and BM3D [Dabov2007BM3D].
+More recently, deep learning approaches, including convolutional neural networks [Jin2017cnn, Guo2019real, wei2020physics, zheng2021adaptive, Fang2025SRENetSL, chi2020dynamic, Elgendy2021LowLightDA], transformers [Liang2021swinir, Lu2022low, zhang2024xformer, Yang2025LearningTS], and diffusion-based models [kawar2022denoising, zhu2023DiffPIR, Nguyen2024diffusion, CheukYiuChan2024AnlightenDiffAD, Wu2025MutuallyRL, Prateek2025quantadiff], have emerged as promising solutions for low-light denoising.
 Low-light image enhancement (LLIE) aims to improve the brightness and contrast of images captured in poorly lit conditions.
-Traditional approaches [16, 31, 29] often leverage heuristic or physical models to enhance image quality, while learning-based methods [28, 56, 24, 5, 17, 37, 53, 59, 22] have demonstrated greater effectiveness when supported by large-scale datasets.
-Recent advancements [45, 38, 42, 58] in this field have increasingly focused on jointly addressing denoising and enhancement tasks, aiming to improve image quality while effectively suppressing noise to maintain natural appearance.
+Traditional approaches [Fu2016Weighted, Li2018Structure, Lee2013Contrast] often leverage heuristic or physical models to enhance image quality, while learning-based methods [Land1971Retinex, Chen2018Retinex, Jiang2021EnlightenGAN, Chen2018See, Guo2020Zero, Lu2021Tbefn, wang2023exposurediffusion, xu2022snr, jiang2023low] have demonstrated greater effectiveness when supported by large-scale datasets.
+Some works also adopt training-free (zero-shot) learning for LLIE [Lv2024FourierPD, Huang2025ZeroShotLI, Xie2024ResidualQL].
+Recent advancements [Shi2024ZERO, Lu2022low, Nguyen2024diffusion, xu2020learning] in this field have increasingly focused on jointly addressing denoising and enhancement tasks, aiming to improve image quality while effectively suppressing noise to maintain natural appearance.
 However, to the best of our knowledge, no specific approach currently exists for joint enhancement and denoising of facial images in low-light conditions.
-Figure 3: Path to the idea. Using facial physical buffers and gallery images can help preserve human face identity information.
 Deep Face Restoration.
-CNN/Transformer-based face restoration methods [54, 14, 23], such as RestoreFormer [54], typically rely on paired low- and high-quality image datasets to learn resolution enhancement and noise reduction.
+CNN/Transformer-based face restoration methods [wang2022RestoreFormer, Fan2022faceformer, Jiang2022dual], such as RestoreFormer [wang2022RestoreFormer], typically rely on paired low- and high-quality image datasets to learn resolution enhancement and noise reduction.
+CodeFormer [zhou2022towards] formulates face restoration as a code prediction task in a discrete latent space, leveraging a Transformer to recover identity-consistent and perceptually realistic results.
 While effective, these methods often struggle under severe degradation and lack the capability to synthesize realistic details in complex scenarios.
 In recent years, generative priors have been increasingly utilized in face restoration to better address real-world degradations.
-GAN-based methods, such as GFP-GAN [52], leverage StyleGAN2 [49] to enhance facial details, while diffusion-based models [55, 7, 13, 48, 51] have further expanded restoration capabilities.
-For instance, BFRDiffusion [7] uses Stable Diffusion to enhance low-quality images by adding high-fidelity details, and Ding et al.
-[13] employ a pretrained diffusion model that denoises degraded inputs while preserving identity by fine-tuning on a person’s gallery photos.
-MGBFR [48] combines text prompts and reference images with a dual-control adapter to retain accurate facial attributes and identity.
-Although these generative methods are effective, they often rely on multi-stage training or finetunning, which can limit the efficient use in some real-world applications.
+GAN-based methods, such as GFP-GAN [wang2021towards], leverage StyleGAN2 [tov2021stylegan2] to enhance facial details, while diffusion-based models [wang2023dr2, Chen2024face, ding2024restoration, Tao2024face, Varanka2024gallery] have further expanded restoration capabilities.
+For instance, BFRDiffusion [Chen2024face] uses Stable Diffusion to enhance low-quality images by adding high-fidelity details, and Ding et al.
+[ding2024restoration] employ a pretrained diffusion model that denoises degraded inputs while preserving identity by fine-tuning on a person’s gallery photos.
+MGBFR [Tao2024face] combines text prompts and reference images with a dual-control adapter to retain accurate facial attributes and identity.
+Although these generative methods are effective, they often rely on multi-stage training or fine-tuning and often struggle recovering identity-consistent face under severe degradation, which can limit the efficient use in some real-world applications.
 Reference Prior for Face Restoration.
-In face restoration, reference images [48, 13, 32, 33] provide essential priors for recovering detailed and high-frequency facial features.
+In face restoration, reference images [Tao2024face, ding2024restoration, li2022learning] provide essential priors for recovering detailed and high-frequency facial features.
 These priors guide the model to generate realistic restorations.
 Reference-based methods can be divided into two types: single reference and gallery-based images.
-The single reference approach [48] uses one image to guide the restoration, focusing on enhancing facial details from a specific reference.
-In contrast, the gallery-based approach [13] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration.
-Despite the effectiveness [13] using gallery photos of users, it still requires multi-stage training and fine-tuning on each user during inference.
-3 Method
-Figure 4: For images affected by mild low-light noise, phycial buffers can be accurately extracted.
-However, when encountering the images degraded by severe low light and noise, precise physical buffers are difficult to obtained.
-3.1 Restoration by Diffusion
-A diffusion model approximates the clean image distribution $p_{\theta}(\mathbf{x}_{0})$ by learning a model $\theta$ that reverses the noise-adding process.
-In Denoising Diffusion Probabilistic Models (DDPMs) [21], Gaussian noise is progressively added to a clean image $\mathbf{x}_{0}$:
-$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\text{where}\quad\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (2)
-The reverse process then removes the noise to reconstruct the clean image.
-After training, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, the model can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t})$.
-In restoration tasks, like low-light denoising, the goal is to recover a high-quality image $\mathbf{x}_{0}$ from a low-quality input $\mathbf{y}$.
-Unlike pure generation, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $\mathbf{y}$.
-During training, the model uses $\mathbf{y}$ as a condition, and once the model is trained, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, it can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y})$ using the trained model.
-Figure 5: The overall architecture of the proposed method.
-Our core idea is to use ID-consistent physical buffers, extracted from gallery photos, to constrain the generative space in the diffusion model restoration process.
-For a high-quality gallery, we use LAP [65] to extract the albedo and normal information for each photo and apply adaptive aggregation to fuse the entire gallery.
-The extracted albedo represents base skin color and facial appearance, while normal captures facial geometry.
-In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
-3.2 Why Combining Gallery and Physical Buffer?
-We formalize our core approach of extracting identity-consistent physical buffers from gallery photos to guide the diffusion model for restoration.
-This method is based on the observation that using facial physical buffers or gallery images can help preserve identity information during restoration tasks involving diffusion models.
-To evaluate the capability of physical buffers and gallery images, we conducted four simple experiments under a blind scheme (excluding $y_{0}$ during training), as illustrated in Fig.
-3:
-  • Exp 1: Model trained solely on generic images (purely generative).
-  • Exp 2: Model trained on generic images, conditioned on the facial physical buffer extracted from the input image.
-  • Exp 3: Model trained on generic images, then fine-tuned with the user’s gallery photos.
-  • Exp 4: Model trained on generic images, fine-tuned with the user’s gallery photos, and conditioned on the facial physical buffer.
-We tested these models on a mildly noisy input.
-Exp 1 produces random faces, as expected from a purely generative model.
-In contrast, Exp 2 uses physical buffers obtained via DECA [15], which can extract coarse face buffers from a single 2D photo (see Fig.
-4), yielding faces with user-like features (e.g., nose shape, mouth and skin tone).
-Hence, physical buffers can convey key identity cues.
-Exp 3 outputs distinct faces of the same user, compared to Exp1, showing that fine-tuning with gallery photos effectively transfers identity information.
-Finally, Exp 4 yields the best results, showcasing that the collaborative use of gallery photos and facial physical prior can achieve the optimal outcomes.
-While Fig. 3 shows us a design direction, we need to overcome:
-  1. (i)
-    Fine-tuning: Extracting identity via fine-tuning requires significant time and must be repeated for each user.
-Additionally, the small size of a gallery compared with a large pre-trained model can lead to overfitting.
-  2. (ii)
-    Unreliable buffers from degraded inputs: As studied in the face reconstruction literature, low-quality inputs consistently present additional challenges [15].
-Severe degradation undermines the extraction of physical buffers (Fig.
-4), resulting in poor reconstruction.
-Moreover, features captured from a single degraded photo are limited, not to mention from a degraded one.
-In conclusion, to avoid repeated fine-tuning on galleries and reduce the risk of erroneous buffers from noisy inputs, it is more desirable to jointly employing gallery photos and facial physical priors, as we will discuss below.
-3.3 Extracting Physical Buffer from Photo Gallery
-Based on the preceding analysis, we propose extracting physical buffers directly from existing clean gallery photos and using them as conditions to guide the diffusion model during training.
-This approach offers significant benefits: clean gallery photos allow us to form comprehensive and robust identity-consistent physical buffers containing more acurate facial identity features of the target person.
-Specifically, We will extract the albedo and normal maps of the target person, which are two key attributes used to describe the physical characteristics of a face.
+The single reference approach [Tao2024face] uses one image to guide the restoration, focusing on enhancing facial details from a specific reference.
+In contrast, the gallery-based approach [ding2024restoration] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration.
+Despite the effectiveness [ding2024restoration] using gallery photos of users, it still requires multi-stage training and fine-tuning on each user during inference.
+The summary comparison between our method and previous methods is summarized in Table I, and the preview of the visual results can be found in Fig. 2.
+III Method
+Our Proposed method is a generative approach based on diffusion models (Section III-A).
+However, a naive diffusion model does not support a personalized gallery as prior.
+The core innovation of our proposed method is the introduction of a physical buffer that can be extracted from the personalized gallery (Section III-B and Section III-C), and then integrated into the diffusion models to achieve the target denoising (Section III-D).
+III-A Denoising Diffusion Model
+Diffusion models [SohlDickstein2015DeepUL, ho2020ddpm] approximate a complex data distribution $q(\mathbf{x}_{0})$ by learning to reverse a gradual noising process that transforms the data into a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$.
+Through this iterative denoising process, they have become the state of the art in generative modeling, outperforming GANs and VAEs in stability and visual fidelity.
+Forward diffusion.
+Given a clean image $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$, the forward process progressively corrupts it with Gaussian noise of increasing variance $\{\beta_{t}\}_{t=1}^{T}$, forming a Markov chain:
+$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}\!\big(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\,\beta_{t}\mathbf{I}\big).$ (2)
+After $T$ steps, the image becomes nearly indistinguishable from Gaussian noise. Using properties of Gaussian composition, the sample at an arbitrary step can be expressed as
+$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I}),$ (3)
+where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ controls the remaining signal-to-noise ratio.
+When $\bar{\alpha}_{t}$ approaches zero, the original signal is nearly destroyed.
+Typical choices of the variance schedule include the linear schedule of Ho et al.
+[ho2020ddpm, dhariwal2021diffusion] and the cosine schedule of Nichol et al.
+[Nichol2021ImprovedDD], which better distribute the noise levels across timesteps and improve sample quality.
+TABLE I: Compared to previous related methods, our approach performs joint low-light enhancement and denoising (low-light), effectively recovering identity-consistent faces even under severe degradation (ID-consistent).
+By leveraging the user’s gallery, our method supports personalized restoration (Personalized), and is scalable to multiple users without requiring fine-tuning when switching users (No fine-tune).
+Method low-light ID-consistent Personalized No fine-tune
+DiffLL [jiang2023low]
+FourierDiff [Lv2024FourierPD]
+MIRNet [zamir2020learning]
+SNR-aware [xu2022snr]
+GFPGAN [wang2021towards]
+CodeFormer [zhou2022towards]
+Ding et al. [ding2024restoration]
+PFStorer [Varanka2024gallery]
+DiffPGD (Ours)
+Reverse denoising.
+The generative (reverse) process learns to invert the corruption by gradually removing noise from $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$ to recover $\mathbf{x}_{0}$.
+A neural network $\epsilon_{\theta}$ is trained to predict the noise component added at each step, thereby parameterizing the reverse conditional
+$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\!\big(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t,\mathbf{c}),\sigma_{t}^{2}\mathbf{I}\big),$ (4)
+where $\mathbf{c}$ denotes any optional conditioning signal. The network is optimized via the simplified denoising objective [ho2020ddpm, rombach2022high] derived from the variational lower bound:
+$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},\,\mathbf{c},\,\epsilon,\,t}\big[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\|_{2}^{2}\big],$ (5)
+where $t$ is uniformly sampled from $\{1,\ldots,T\}$. Intuitively, $\epsilon_{\theta}$ learns a vector field that points from noisy samples back to the data manifold at different noise scales.
+Sampling process. At inference, given a noisy image $\mathbf{x}_{t}$, the model iteratively refines it by sampling from the learned reverse process:
+$\mathbf{x}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\right)+\sigma_{t}z,$ (6)
+where $z\sim\mathcal{N}(0,\mathbf{I})$ and typically $\sigma_{t}^{2}=\beta_{t}$. This iterative denoising gradually reconstructs the clean data $\mathbf{x}_{0}$ from pure noise $\mathbf{x}_{T}$.
+This forward–reverse framework provides a simple yet powerful formulation that unifies generative modeling, denoising, and Bayesian inference.
+By adjusting the conditioning signal $\mathbf{c}$, the same architecture can handle diverse tasks such as image restoration, super-resolution, and text-guided generation.
+In restoration settings, $\mathbf{c}$ often represents the degraded observation or auxiliary features (e.g., edge, illumination, or identity), and the reverse process recovers a clean image consistent with this condition.
+Figure 2: Preview of the visual results.
+Using gallery photos from a user’s smartphone, we can restore low-light, noisy facial images.
+Our method produces finer details and better identity compared to state-of-the-art low-light and face restoration approaches like MIRNet [zamir2020learning], DiffLL [jiang2023low], GFP-GAN [wang2021towards], FourierDiff [Lv2024FourierPD], and CodeFormer [zhou2022towards].
+Please zoom in for the best visual experience.
+III-B Why Combining Gallery and Physical Buffer?
+Motivation.
+After introducing the diffusion framework in Section III-A, we recall that a conditional diffusion model requires additional information $\mathbf{c}$ to constrain the reverse process.
+For personalized restoration, such a condition must encode the subject’s identity to prevent hallucination.
+We therefore investigate two possible priors for identity preservation: (1) the gallery prior, and (2) the physical-buffer prior.
+A gallery refers to a small set (typically 3–6) of high-quality face photos of the target person, captured under different poses and lighting conditions.
+These images are readily available from user photo collections.
+They contain rich identity cues, but are also affected by background, expression, and illumination variations.
+A physical buffer refers to an intermediate facial representation—in our case the albedo and surface normal—that describes the physical properties of a person’s face independent of lighting and shading.
 Albedo captures the intrinsic color of an object’s surface—that is, the subject’s color properties without any lighting effects.
@@
 eyebrows, lip coloration, and other distinct texture elements can reflect individual identity traits.
-  • Normal encodes the surface orientation at each pixel, which provides hints about the underlying shape characteristics.
-By conditioning on these identity physical buffers, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction.
+Surface normal encodes the surface orientation at each pixel, which provides hints about the underlying shape characteristics.
+These buffers can be extracted using pretrained face reconstruction networks [feng2021learning, Zhang2021lap].
+Compared to the gallery, they provide disentangled, lighting-independent identity features, but are less directly visual.
+Controlled Comparison. To analyze the complementary effects of these priors, we conduct four controlled generative experiments (Fig. 3).
+  • Experiment 1: Model trained solely on generic images from the FFHQ dataset [Karras2018ASG].
+  • Experiment 2: Model trained on generic images, while conditioned on the facial physical buffer extracted from the noisy input image.
+  • Experiment 3: Model trained on generic images, then fine-tuned with the user’s gallery photos.
+  • Experiment 4: Model trained on generic images, fine-tuned with the user’s gallery photos, and conditioned on the facial physical buffer extracted from the input image.
+We tested these trained models on a mildly low-light noisy photo of a user, and the results are presented in the Fig.
+3 (right).
+The model from Exp 1 produces random faces, as expected from a purely generative purpose.
+In contrast, the model from Exp 2 uses some coarse face physical buffers extracted from the input image [feng2021learning] as condition.
+The result from Exp 2 yields faces with user-like features (e.g., nose shape, mouth shape and skin tone).
+Hence, even just some coarse physical buffers extracted from a degraded photo can convey some key identity cues.
+Exp 3 outputs distinct faces of the same user, compared to Exp 1, showing that fine-tuning with gallery photos transfers identity information.
+Finally, Exp 4 yields the best results, showcasing that the collaborative use of gallery photos and facial physical prior can achieve the optimal outcomes.
+Figure 3: Incorporating gallery photos and physical buffers enhances identity preservation.
+Exp 1: Trained on generic images only $\rightarrow$ outputs a generic face.
+Exp 2: Adds physical buffer conditioning $\rightarrow$ captures some user-specific features.
+Exp 3: Fine-tuned with gallery photos $\rightarrow$ generates an arbitrary user face.
+Exp 4: Combines gallery fine-tuning and physical buffer $\rightarrow$ restores the expected user face.
+III-C Extracting Physical Buffer from Photo Gallery
+While Fig. 3 shows us a design direction, there are still two fundamental limitations we need to overcome:
+  1. (i)
+    Fine-tuning: Extracting identity via fine-tuning requires significant time and must be repeated for each user [ding2024restoration].
+Additionally, the small size of a gallery compared with a large pre-trained model can lead to overfitting.
+  2. (ii)
+    Unreliable buffers from degraded inputs: As studied in the face reconstruction literature, low-quality inputs consistently present additional challenges [feng2021learning].
+Fig.
+4 shows that extracting albedo and normal maps directly from degraded inputs becomes unreliable as the noise level increases, often leading to incorrect physical buffers, resulting in poor reconstruction.
+Moreover, features captured from a single image are coarse and limited, not to mention from a degraded one.
+Our approach: Extract Physical Buffer from the gallery.
+Based on the preceding analysis, we propose extracting physical buffers directly from the clean gallery photos instead of noisy input image.
+This approach offers significant benefits: The gallery provides diverse but unstructured identity evidence, while the physical buffer offers structured, lighting-independent representations.
+By deriving the physical buffers from the clean gallery, we combine these complementary strengths—retaining the gallery’s diversity while achieving physically consistent and accurate facial identity features of the target person.
+The resulting ID-consistent buffers remain reliable even under severe noise and serve as a strong condition for the diffusion-based restoration process.
+By conditioning on the identity physical buffers directly from existing clean gallery photos, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction.
 Consequently, the model is better equipped to preserve the target person’s identity and facial features, even under severe degradation.
-However, extracting physical buffers for a person from a set of gallery photos is a more challenging task than extracting these buffers from a single input, as in methods like DECA.
-This complexity arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses.
+Figure 4: For input images affected by mild low-light noise, physical buffers can be accurately extracted from input.
+However, when encountering the input images degraded by severe low light and noise, precise physical buffers are difficult to obtain from input.
+How to extract physical buffer from gallery?
+However, extracting physical buffers from a set of gallery photos is more challenging than extracting from a single input.
+The complication arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses.
 To ensure these buffers comprehensively represent the person’s identity and facial properties, they must be aligned to maintain identity consistency.
-As shown in the physical buffer extractor module in Fig.
-5(a), we adopt the aggregation net design from the LAP [65] to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
-3.4 Model Framework
-We name our final proposed model DiffPGD.
+Inspired by LAP [Zhang2021lap], we used an aggregation network as the physical buffer extractor module (shown in Fig.
+5, to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
+Details of the proposed scheme.
 As illustrated in Fig.
-5, the model leverages ID physical buffers (albedo $a^{i}$ and normal $n^{i}$) extracted from person i’s gallery photos as conditions to guide our diffusion model during training directs the generative process toward a personalized space.
+5, the physical buffer extractor gets the ID physical buffers (albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$) extracted from person i’s gallery photos.
 The aggregation network contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting the consistent representation.
-To model albedo and normal, two separate aggregation networks, denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed.
-Given a photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i, each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $\mathbf{\alpha}_{k}^{a}$ and $\mathbf{\alpha}_{k}^{n}$, respectively.
-To derive a global representation of the identity for person i based on $\{\mathbf{\alpha}_{k}^{a},\mathbf{\alpha}_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted, This method learns a weight set $\{\mathbf{w}_{k}^{a},\mathbf{w}_{k}^{n}\}_{k=1}^{N}$ according to the quality of each image in$\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$.
-The combined global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ for texture and depth are calculated as:
-$\mathbf{\alpha}_{c}^{a}=\sum_{k=1}^{N}\mathbf{w}_{k}^{a}\mathbf{\alpha}_{k}^{a},\text{ and}\quad\mathbf{\alpha}_{c}^{n}=\sum_{k=1}^{N}\mathbf{w}_{k}^{n}\mathbf{\alpha}_{k}^{n}.$ (3)
-Once calculated, the global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ are passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$.
-During inference, the process begins with a Gaussian noise image $\mathbf{x}_{T}^{i}$, which is iteratively transformed into the clean, enhanced image $\mathbf{x}_{0}^{i}$ by sampling from $p(\mathbf{x}_{0}^{i}|\mathbf{x}_{T}^{i},\mathbf{y}^{i},\mathbf{a}^{i},\mathbf{n}^{i})$ using the trained model.
-With the low-light noisy image $\mathbf{y}^{i}$ as another condition during training, other features, such as pose, hairstyle, and background inside the image beyond facial features can be effectively captured as well.
+To model albedo and normal, two separate aggregation networks, denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed (Fig.
+5 uses the albedo extraction as example, and the normal extraction follows the similar strategy).
+Given a photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i (we will omit the identity index i in the following for simplicity), each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $\mathbf{\alpha}_{k}^{a}$, $\mathbf{\alpha}_{k}^{n}$ $\in$ $[1,1,c]$, respectively.
+Due to the quality of $\{\textbf{F}_{k}\}_{k=1}^{N}$, the importance of each dimension in $\mathbf{\alpha}_{k}^{a}$, $\mathbf{\alpha}_{k}^{n}$ which reveals how correlated it is to the identity, is supposed to be different.
+Therefore, to derive a global representation of the identity based on $\{\mathbf{\alpha}_{k}^{a},\mathbf{\alpha}_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted.
+More specifically, channel-wise weights $\mathbf{w}_{k}^{a},\mathbf{w}_{k}^{n}$ $\in$ $[1,1,c]$ are also learned to represent the importance of each dimension in $\mathbf{\alpha}_{k}^{a}$, $\mathbf{\alpha}_{k}^{n}$:
+[…]
+A soft-max function is further used to normalize the weights into $\{\hat{\mathbf{w}}_{k}^{a}\}_{k=1}^{N},\{\hat{\mathbf{w}}_{k}^{n}\}_{k=1}^{N}$.
+Finally, the combined global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ for texture and depth are calculated as:
+$\mathbf{\alpha}_{c}^{a}=\sum_{k=1}^{N}\mathbf{\hat{w}}_{k}^{a}\cdot\mathbf{\alpha}_{k}^{a},\text{ and}\quad\mathbf{\alpha}_{c}^{n}=\sum_{k=1}^{N}\mathbf{\hat{w}}_{k}^{n}\cdot\mathbf{\alpha}_{k}^{n}.$ (8)
+Once calculated, the global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ are passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $\mathbf{a}$ and normal $\mathbf{n}$:
+$\mathbf{a}=\phi^{a}(\mathbf{\alpha}_{c}^{a}),\text{ and}\quad\mathbf{n}=\phi^{n}(\mathbf{\alpha}_{c}^{n}).$ (9)
+The algorithm for our physical buffers extraction process is shown in Algorithm 1.
+Figure 5: ID-consistent physical buffer extractor.
+The extractor gets the ID physical buffers extracted from target person’s gallery photos.
+It contains a shared encoder across multiple images and a global decoder for predicting the consistent representation.
+Note we only show albedo here, and the normal extractor follows the similar strategy.
+The module weights are initialized from the aggregation network in [Zhang2021lap] and frozen during training.
+Algorithm 1 Extracting physical buffers from photo gallery
+1:Input: A photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i (we omit the identity index i in the following for simplicity),
+[…]
+[…]
+[…]
+[…]
+6:return $\mathbf{a},\mathbf{n}$
+Figure 6: Overview of our method.
+We leverage ID-consistent physical buffers (albedo and normal) extracted from gallery photos to constrain the generative space during diffusion restoration.
+The albedo encodes intrinsic skin/appearance while the normal represents facial geometry.
+In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
+III-D Diffusion-based Restoration Model Framework
+In Section III-C we discussed how to extract the physical buffer.
+Now, we discuss how this physical buffer can be used in the restoration pipeline.
+We name our final proposed model DiffPGD, which adopts a diffusion-based restoration framework.
+As illustrated in Fig.
+6, the model leverages ID physical buffers (albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$) extracted from person i’s gallery photos as conditions to guide our diffusion model during training to direct the generative process toward a personalized space.
+We will drop the superscript $i$ in the following for simplicity.
+In restoration tasks, such as low-light denoising task solved in this paper, the goal is to recover a high-quality image $\mathbf{x}_{0}$ from a low-quality input $\mathbf{y}$.
+Unlike a pure generation task, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $\mathbf{y}$.
+As shown in the conditional denoising diffusion model in Fig.
+6, the forward/diffusion process (from right to left) gradually adds Gaussian noise, denoted by $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$.
+Our goal is to reverse the diffusion process (from left to right) by gradually recovering the image from the input Gaussian noise with conditions from different modalities, which corresponds to learning the reverse process of a fixed Markov Chain of length $T$ conditioned on $\mathbf{y}$ and the facial physical buffers albedo and normal $(\mathbf{a},\mathbf{n})$, extracted from the photo gallery.
+When training our model, we adopt a P2-weighted loss format [choi2022perception] with an additional weight term $\lambda_{t}$, and the corresponding objective function measures the discrepancy between the predicted and ground-truth noise, and can be formatted as:
+$\mathcal{L}=\lambda_{t}\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},\mathbf{a},\mathbf{n},\epsilon,t}[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{y},\mathbf{a},\mathbf{n})\|_{2}^{2}],$ (10)
+According to Eq.
+(10), $\epsilon_{\theta}$ takes as input the noisy target image $\mathbf{x}_{t}$, the time step $t$, and the conditional signal $\mathbf{c}$, which consists of the low-quality input image $\mathbf{y}$ and the extracted physical buffers $(\mathbf{a},\mathbf{n})$, to estimate the noise $\epsilon$.
+$\lambda_{t}$ serves as a hyperparameter to adjust the loss weight dynamically across different time step.
+It assigns higher weights to the loss at levels where the model learns perceptually rich contents while minimal weights to which the model learns imperceptible details.
+During inference, the process begins with a pure Gaussian noise image $\mathbf{x}_{T}\sim\mathcal{N}(\textbf{{0}},\textbf{{I}})$, the model learns the conditional transition distribution $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y},\mathbf{a},\mathbf{n})$ and iteratively denoise the image for $T$ steps, generating the target image $\mathbf{x}_{0}$ in the end.
+Following the notations defined in Sec.
+III-A, the training and inference processes of our model is summarized in Algorithm 2 and 3.
+Algorithm 2 Training our denoising model $\epsilon_{\theta}$
+[…]
+2:repeat
+3:  $(\mathbf{x}_{0},\mathbf{y})\sim q(\mathbf{x}_{0},\mathbf{y})$
+4:  $t\sim\text{Uniform}({1,...,T})$
+5:  $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
+6:  $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$
+7:  Take a gradient descent step on
+$\nabla_{\theta}\lambda_{t}\left\|\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{y},\mathbf{a},\mathbf{n}\right)-\epsilon\right\|_{2}^{2}$
+8:until converged
+Algorithm 3 Inference in $T$ iterative refinement steps
+[…]
+[…]
+3:for $t=T,\ldots,1$ do
+4:  $z\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else $z=0$
+5:  $\mathbf{x}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{y},\mathbf{a},\mathbf{n})\right)+\sigma_{t}z$
+6:end for
+7:return $\mathbf{x}_{0}$
 Our experiments demonstrate that identity information is effectively preserved with the guidance of these ID-consistent buffers, enabling our model to achieve superior results in human face restoration.
-4 Experiments
-Figure 6: Visual comparison on real cases.
-The first column presents the input images along with their capture parameters.
-In the second column, we enhance the brightness to better reveal the noise.
-Please zoom in for the best visual experience.
-4.1 Dataset
+IV Experiments
+IV-A Dataset
 Training Data.
-Since no publicly available datasets exist for low-light noisy face images, we simulate such data using the CelebAMask-HQ dataset [30], the InverseISP algorithm [34], and the Poisson-Gaussian noise model Eq.
-1.
-The pipeline involves converting RGB images to their corresponding RAW mosaiced format using InverseISP [34], followed by applying Eq.
-1 to simulate low-light noise using realistic camera sensor parameters.
-Multiple noise levels are generated by varying the light intensity (Lux) to ensure the model is robust across diverse noise conditions.
-Refer to supplementary for more details on the simulation.
-  Photons Per Pixel 5 10 13 26
-Methods ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$
-SNR-aware[59] 0.2156 188.3311 0.1302 8.4541 0.4240 139.8193 0.0786 9.5462 0.4999 124.1162 0.0626 10.1712 0.6937 99.9875 0.0383 12.4961
-MIRNet[61, 62] 0.1229 294.7381 0.2414 8.6008 0.2860 214.7785 0.1534 9.8572 0.3830 187.2571 0.1286 10.4209 0.6212 136.1320 0.0793 12.8992
-DiffLL[22] 0.2910 167.3905 0.1068 17.3506 0.4544 150.4559 0.1117 18.3360 0.5095 140.2738 0.0996 19.9547 0.6833 119.2693 0.0760 23.3180
-GFPGAN[52] 0.1941 388.9293 0.4256 7.1990 0.2075 233.3173 0.1521 8.1283 0.2976 190.7264 0.1326 9.5483 0.5976 158.9475 0.0829 11.4855
-CodeFormer[67] 0.2863 111.4544 0.04908 18.2406 0.4940 91.1244 0.0326 19.2692 0.5563 84.7955 0.0281 20.2306 0.7338 80.6582 0.0263 24.5512
-Ours 0.3714 77.5743 0.0084 17.9334 0.5225 71.8645 0.0063 18.4999 0.5938 69.5567 0.0042 19.0910 0.7118 59.1975 0.0030 23.7226
-Table 1: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset on 5PPP, 10PPP, 13PPP, and 26PPP. The lower the PPP, severe the degradation is.
-Testing Data. For testing DiffPGD, we utilize both simulated and real-world captured low-light noisy data.
-The simulation setup assumes a $12$-bit image sensor with parameters similar to that of Canon DSLR cameras: a Full Well Capacity of $36000\text{e}^{-}$, Quantum Efficiency of $0.42$, dark current of $7\text{e}^{-}$, and read noise of $6\text{e}^{-}$.
+Since no publicly available datasets exist for low-light noisy face images, we simulate such data using the CelebAMask-HQ dataset [Lee2020CelebAMaskhq], the InverseISP algorithm [li2024efficient], and the Poisson-Gaussian noise model (Eq.
+1).
+As shown in Fig.
+7, the first step in the simulation process involves passing the RGB image through a pre-trained Inverse ISP network [li2024efficient] to generate a pseudo ground truth RAW mosaic image.
+For visualization purposes, we display the demosaiced version of this image.
+We utilize the pretrained weights provided [li2024efficient], which were obtained by training the network on iPhone camera images for unprocessing the RGB images into RAW format.
+The pseudo ground truth RAW image then passed through a Poisson-Gaussian simulator to simulate the low-light noisy RAW mosaic image.
+During the training process, we vary the camera sensor parameters across a wide range of values to allow the model to learn diverse camera sensor noise profiles.
+The simulation setup assumes a $12$-bit image sensor with parameters similar to typical DSLR cameras.
+The parameter ranges are summarized in Table V.
+TABLE V: Simulation parameters used in the Poisson-Gaussian Simulator.
+  Parameter Range
+Full Well Capacity (FW) $19000\text{e}^{-}\leq\text{FW}\leq 64000\text{e}^{-}$
+Quantum Efficiency ($\mathrm{\eta}$) $0.32\leq\mathrm{\eta}\leq 0.54$
+Dark Current ($\mathrm{\sigma_{dark}}$) $2.2\text{e}^{-}\leq\mathrm{\sigma_{dark}}\leq 11.7\text{e}^{-}$
+Read Noise ($\mathrm{\sigma_{read}}$) $2.2\text{e}^{-}\leq\mathrm{\sigma_{read}}\leq 10.8\text{e}^{-}$
+photo level ($\mathrm{\alpha}$) $13\,\text{ppp}\leq\mathrm{\alpha}\leq 65\text{ppp}$
+Testing Data. To test DiffPGD, we utilize both simulated and real-world captured low-light noisy data.
+For the simulated testing data, similar to the training data simulation, the simulation setup assumes a $12$-bit image sensor with parameters similar to that of a DSLR camera: a full well capacity (FW) of $36000\text{e}^{-}$, quantum efficiency ($\mathrm{\eta}$) of $0.42$, dark current of $\mathrm{\sigma_{dark}}=7\text{e}^{-}$, and read noise of $\mathrm{\sigma_{read}}=6\text{e}^{-}$.
 These parameters ensure realistic noise characteristics for evaluating model performance.
@@
 We cropped these images into 256×256 face patches.
-4.2 Implementation Details
-Our denoising model is built on the ADM architecture [12], and the ID-consistent physical extractor is implemented using the aggregation network from LAP [65].
-For training, we use a P2-weighted loss function [8], which measures the discrepancy between the predicted and ground-truth noise.
-This loss is defined as: $L=\lambda_{t}||\hat{\epsilon}_{t}-\epsilon_{t}||_{2}^{2}$ , where $\lambda_{t}$ serves as a hyperparameter to adjust the loss weight dynamically across different timesteps.
-During the training stage, we train our model on the CelebAMask-HQ dataset [30], comprising 30,000 images distributed across 4,516 unique identities(IDs).
+Figure 7: Simulation process for generating low-light noisy face images.
+A ground truth RGB image is first passed through a pre-trained Inverse ISP [li2024efficient] to obtain a pseudo ground truth RAW mosaic.
+This is followed by applying a Poisson-Gaussian noise model with varying sensor parameters to simulate a low-light noisy RAW.
+Finally, a Forward ISP module converts the noisy RAW back to a low-light noisy RGB image.
+IV-B Implementation Details
+Our denoising model is built upon the ADM architecture [dhariwal2021diffusion].
+As illustrated in Fig.
+6, we modify the original architecture so that the model can take concatenated degraded images as conditional inputs.
+In addition, we employ an encoder (ResNet-18 [he2016deep]) to extract a global latent code from the target individual’s physical buffers, which are obtained from the gallery photos.
+Following the FiLM-style [Perez2017FiLMVR] modulation used for time embeddings in ADM [dhariwal2021diffusion], this latent code modulates the UNet features in each residual block via scale–and–shift or additive transforms, thereby injecting identity information into the diffusion model.
+The model parameters are configured as follows: the number of channels is set to 128, channel multiple is (1, 1, 2, 2, 4, 4), head channels are 128, attention resolution is 16, dropout rate is 0.1, diffusion steps are 1000.
+For additional architectural details, refer to [dhariwal2021diffusion].
+The ID-consistent physical extractor is directly initialized from the pre-trained weights of the aggregation network in LAP [Zhang2021lap], without the need of retraining.
+For training our model, we use the loss function defined in Eq.
+(10).
+The image size is 256 $\times$ 256.
+During the training stage, we train our model on the CelebAMask-HQ dataset [Lee2020CelebAMaskhq], comprising 30,000 images distributed across 4,516 unique identities (IDs).
 We use 3318 IDs for training and 200 IDs for simulated testing.
 Within each ID, there are 3–20 images.
-The number of gallery images we used per person varies from 3 to 6, with most around 5.
-The Adam optimizer is applied with a learning rate of 1e-4.
+The number of gallery images we used per person varies from 3 to 6, with most of them around 5.
+During inference, we used 3 to 6 images per person for the simulated test set, and 6 gallery images per person for the real-captured test set.
+The model is trained using the Adam optimizer [kingma2014adam] with a learning rate of $10^{-4}$.
+The weight decay is set to 0.
 Training is performed for 200,000 iterations with a batch size of 32, resulting in a total of 6,400,000 samples seen by the model.
-4.3 Performance in Real-Captured Scenarios
+During inference, we adopt the DDIM method [song2020denoising] with 200 denoising steps.
+All experiments are performed on a single NVIDIA A100 GPU.
+IV-C Performance in Real-Captured Scenarios
 Testing model performance in real-life low-light scenarios is crucial and challenging, as real-world conditions often involve uncontrollable factors compared to simulated ones.
-Since no existing data set addresses the restoration of real-life low-light noisy face photos, we captured photos of real persons under indoor and outdoor low-light scenarios , as illustrated in Section 4.1.
-We compare our methods with several state-of-the-art works: UTV-Net[66], GFPGAN[52], SNR-aware[59], MIRNet[61, 62], DiffLL[22], Codeformer [67].
+Since no existing data set addresses the restoration of real-life low-light noisy face photos, we captured photos of real persons under indoor and outdoor low-light scenarios, our procedure is described in Section IV-A.
+We compare our method with several state-of-the-art methods: UTV-Net [zheng2021adaptive], GFPGAN[wang2021towards], SNR-aware[xu2022snr], MIRNet[zamir2020learning], DiffLL[jiang2023low], Codeformer [zhou2022towards].
 These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research.
-Particularly, GFPGAN and Codeformer are designed specifically for face image restoration, which is highly competitive with our method.
+Particularly, GFPGAN[wang2021towards] and Codeformer[zhou2022towards] are designed specifically for face image restoration, which is highly competitive with our method.
 We re-trained/fine-tuned all the baselines with the same training dataset as ours for fair comparison.
 Quantitative and qualitative results are presented.
-To assess the quality of restored images, we employ the Fréchet Inception Distance (FID) [20] and the Kernel Inception Distance (KID) [1], as these metrics offer a robust evaluation aligned with human perception.
-Since ground-truth images are unavailable for real-captured cases, we compute the FID and KID score between the model’s output and the corresponding subjects’ gallery photos.
-As shown in Fig.
-6, our method produces images with the highest quality and level of detail.
-The restored images exhibit more natural skin tones and well-preserved facial identity features, such as eye shape, eyebrows, and lips, while maintaining the highest consistency in facial identity.
+To assess the quality of restored images, we employ the Fréchet Inception Distance (FID) [Heusel2017FID] and the Kernel Inception Distance (KID) [Binkowski2018KID], as these metrics offer a robust evaluation aligned with human perception.
+Since ground-truth images are unavailable for real-captured cases, we compute FID and KID scores between the model’s output and the corresponding subjects’ gallery photos.
+The results are shown in Table VI.
+Our method achieves the best FID score and the second-best KID score.
+When examining the visual results presented in Figs.
+8, 9 and 10, our method produces images with the highest perceptual quality and level of detail.
+The restored images exhibit more natural skin tones and well-preserved facial identity features, such as eye shape, eyebrows, and lips.
+Figure 8: Visual comparison on real cases.
+The first image presents the input images along with their capture parameters.
+In the second image, we enhance the brightness to better visualize the noise present in the input.
+Please zoom in for the best visual experience.
+Figure 9: Visual comparison on real cases.
+The first image presents the input images along with their capture parameters.
+In the second image, we enhance the brightness to better reveal the noise in the input.
+Please zoom in for the best visual experience.
+Figure 10: Visual comparison on real cases.
+The first image presents the input images along with their capture parameters.
+In the second image, we enhance the brightness to better reveal the noise inside the input.
+Please zoom in for the best visual experience.
+Figure 11: Visual comparison on the simulated cases.
+The lower the ppp, the more severe the degradation is.
+Our method maintains good identity consistency even at extremely low ppp and exhibits fewer artifacts.
+Please zoom in for the best visual experience.
 Since our work focuses on face images, we also check on the identity preservation performance.
-Besides the image quality metrics, we adopt the identity score which calculates the cosine similarity of the image features extracted by ArcFace[11] to further evaluate the identity consistency.
+Besides the image quality metrics, we adopt the identity score, which computes the cosine similarity of face embeddings extracted by the FaceNet model [Schroff2015FaceNetAU] pretrained on the VGGFace2 dataset [Cao2017VGGFace2AD], to evaluate identity consistency.
 Again, since ground truth images are unavailable for real-captured cases, we calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo of the corresponding subjects.
-The results are also shown in Table 2.
-Again, our method has the highest identity score, and the visual results are consistent with the quantitative metric results.
-The results demonstrate the supreme ability of our proposed DiffPGD model on real low light denoising and enhancement task.
-4.4 Performance in Simulated Scenarios
+The results are shown in Table VI.
+Our method has the highest identity score, and the visual results are consistent with the quantitative metric results, maintaining the highest consistency in facial identity.
+The results demonstrate the ability of our proposed DiffPGD model on real low light denoising and enhancement tasks.
+TABLE VI: Comparisons with state-of-the-art works on the real-captured test cases.
+  Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
+SNR-aware[xu2022snr] 358.03 0.3206 0.56
+UTVNet[zheng2021adaptive] 360.13 0.2304 0.42
+MIRNet[zamir2020learning] 274.57 0.1732 0.63
+DiffLL[jiang2023low] 301.13 0.2496 0.69
+GFPGAN[wang2021towards] 276.30 0.2023 0.30
+CodeFormer[zhou2022towards] 295.68 0.2158 0.61
+Ding et al. [ding2024restoration] 295.85 0.2438 0.24
+FourierDiff[Lv2024FourierPD] 303.78 0.2383 0.70
+Ours 270.14 0.1765 0.72
+IV-D Performance in Simulated Scenarios
 In simulated scenarios, we use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set.
 We simulated various low-light noisy levels by setting different photons per pixel (ppp), i.e., the average number of photons arriving at the sensor within the exposure time.
-The lower the PPP, severe the degradation is.
-A lower photon level increases the impact of read noise and dark current on the final output, thereby decreasing the quality of the reconstructed images.
-Quantitative results are provided in Table 1.
-The image restored from our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistent best.
-(Refer to supplementary for more visual comparisons).
-From the table, we could see that CodeFormer surpasses us on PSNR.
+A low photon level increases the impact of read noise and dark current on the final output, thereby decreasing the quality of the reconstructed images.
+Quantitative results are provided in Table VII.
+The image restored by our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistency.
+From the table, we could see that CodeFormer surpasses us on PSNR value.
 Fig.
-7 illustrates a visual comparison between our method and CodeFormer under inputs ranging from 5 to 65 PPP.
-We observe that CodeFormer tends to generate smoother images while we could preserve better high-frenquency details.
-Furthermore, our method demonstrates superior identity preservation even at low resolutions, such as 5 PPP to 13 PPP, while exhibiting fewer artifacts.
+11 illustrates a visual comparison between our method and state-of-the-art works under inputs ranging from 5 to 39 ppp, which corresponds to $\sim$0.049 lux to 0.384 lux, estimated at the sensor side assuming typical low-light CMOS settings with pixel pitch in the range of 2–3 $\mu\text{m}$, quantum efficiency 0.3–0.54, and short integration time on the order of 10–30 ms.
+We observe that CodeFormer tends to generate smoother images while we could preserve better high-frequency details.
+Furthermore, our method demonstrates superior identity preservation even at low resolutions, such as 5 ppp to 13 ppp ($\sim$0.049 lux to 0.128 lux), while exhibiting fewer artifacts.
 The advantage of our model in maintaining identity consistency is particularly evident in heavily degraded cases.
 The results highlight the strong performance of our proposed DiffPGD model in low-light denoising and enhancement tasks.
+Like any denoising method, there is a physical limit beyond which our method will break down.
+This happens when input image is below 5ppp ($\sim$0.049 lux).
+However, in this extremely low photon condition, we still perform much better than competing methods.
+TABLE VII: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset on 5ppp ($\sim$0.049 lux), 10ppp($\sim$0.098 lux), 13ppp($\sim$0.128 lux), 26ppp($\sim$0.256 lux), and 39ppp($\sim$0.384 lux).
+The lower the ppp, the more severe the degradation is.
+  lux 0.049 lux 0.098 lux 0.128 lux 0.256 lux 0.384 lux
+Methods ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$
+SNR-aware[xu2022snr] 0.3093 200.10 0.1382 0.5580 134.44 0.0664 0.6366 120.44 0.0526 0.7922 96.55 0.0330 0.8547 84.36 0.0244
+MIRNet[zamir2020learning] 0.1256 376.98 0.4035 0.3596 320.81 0.2719 0.4698 303.99 0.2506 0.7068 209.01 0.1579 0.7928 161.86 0.1106
+FourierDiff[Lv2024FourierPD] 0.0451 384.58 0.4545 0.2045 308.99 0.2753 0.3233 275.99 0.2121 0.6118 202.49 0.1462 0.7195 163.25 0.1137
+DiffLL[jiang2023low] 0.3835 190.41 0.1687 0.5496 167.77 0.1390 0.6228 155.87 0.1244 0.7796 128.73 0.0898 0.8388 115.59 0.0776
+GFPGAN[wang2021towards] 0.0070 388.93 0.4256 0.1194 233.32 0.1522 0.2851 203.81 0.1152 0.6175 123.27 0.0624 0.7294 106.47 0.0497
+CodeFormer[zhou2022towards] 0.3553 111.46 0.0491 0.5867 91.13 0.0326 0.6501 84.80 0.0281 0.7998 80.66 0.0264 0.8645 64.04 0.0167
+Ours 0.4527 85.50 0.0178 0.6236 75.39 0.0135 0.6923 75.65 0.0179 0.8168 68.44 0.0114 0.8719 62.73 0.0112
+V Ablation Study
+ID-Consistent Physical Buffers for Identity Preservation.
+In Table VIII, we compare the quantitative results of training our model with and without ID physical buffers in real-captured test cases.
+The results demonstrate that using physical buffers significantly increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
+TABLE VIII: Ablation comparisons of using ID physical buffers on real-captured test cases.
   Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
-SNR-aware[55] 252.2397 0.1264 0.1990
-UTVNet[66] 341.8502 0.1617 0.2002
-MIRNet[61, 62] 276.3853 0.1659 0.2210
-DiffLL[22] 252.1553 0.1421 0.3196
-GFPGAN[52] 252.7411 0.1249 0.1525
-CodeFormer[67] 281.9744 0.1459 0.2965
-Ours 251.9318 0.0882 0.3708
-Table 2: Comparisons with state-of-the-art works on the real-captured test cases.
-  Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
-w/o physical buffers 272.6148 0.1281 0.3169
-w/ physical buffers 251.9318 0.0882 0.3708
-Table 3: Ablation comparisons of using ID physical buffers on real-captured photo test.
-Figure 7: Visual comparison on the simulated cases. Our method maintains good identity consistency even at extremely low PPP and exhibits fewer artifacts.
-Figure 8: The relationship between the identity (ID) score and FID of the restored images and the number of gallery images used to generate the ID physical buffer.
-5 Ablation Study
-ID-Consistent Physical Buffers for Identity Preservation.
-We have shown that the proposed model DiffPGD exhibits outstanding performance in both simulated and real-captured testing scenarios.
-In Table 3, we compare the quantitative results of training our model with and without ID physical buffers in real-captured test cases.
-The results demonstrate that using physical buffers largely increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
+w/o physical buffers 281.68 0.1814 0.68
+w/ physical buffers 270.14 0.1765 0.72
 Effect of the Number of Gallery Images.
 The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the user’s gallery photos.
-As claimed in the paper, facial features extracted from a single image are inherently limited.
+As stated in the paper, facial features extracted from a single image are inherently limited.
 In Fig.
-8, we evaluate the effect of the number of gallery images used to generate the ID physical buffer.
+12, we evaluate the effect of the number of gallery images used to generate the ID physical buffer.
 We analyze it with the real-captured set where 6 gallery images were collected from each subject.
 The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image.
-Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it, which corresponds to the case without gallery photos.
-6 Conclusion
+Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it.
+Figure 12: The relationship between the identity (ID) score and FID of the restored images and the number of gallery images used to generate the ID physical buffer.
+Direct-Gallery vs.
+Physical Buffer?
+Some previous studies have attempted to extract latent facial feature representations directly from reference photos [Varanka2024gallery].
+To evaluate whether this strategy can effectively capture sufficient facial information to guide the diffusion model, compared with extraction from physical buffers, we conduct an experiment.
+As illustrated in Fig.
+13, instead of deriving features from physical buffers, we directly extract latent codes from the target individual’s gallery photos.
+Figure 13: A network architecture diagram for identity feature restoration.
+It extracts latent identity features from the user’s photo gallery directly with an encoder, which are then used to condition a denoising diffusion model.
+TABLE IX: Comparison between ID physical-buffer-based and direct-gallery-based methods on the real-captured photo test set.
+For a fair comparison, the same gallery set is used for each individual in both the direct-gallery-based and physical-buffer-based approaches.
+  Methods LPIPS $\downarrow$ SSIM $\uparrow$ KID $\downarrow$ ID Score $\uparrow$
+w/ direct-gallery 0.6316 0.2574 0.0371 0.77
+w/ physical buffers 0.5742 0.4175 0.0161 0.83
+Figure 14: Visual comparison on different identity representation strategies.
+We compare the restored images generated with the direct-gallery-based method with the proposed physical buffer-guided method on our real-captured photo set.
+This is a special semi-real set, in which we take photos of printed paper photos at night, allowing ground-truth reference for evaluation.
+The quantitative and visual comparisons are presented in Table IX and Fig.
+14, respectively.
+The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the direct-gallery-based approach.
+One major reason for this is that the ID physical buffers serve as robust conditions, enabling the extraction of comprehensive facial features from the gallery.
+In contrast, the direct-gallery-based method is less focused on facial features and is often influenced by irrelevant factors such as background and hairstyle.
+Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
+VI Conclusion
 In this work, we addressed the challenges of low-light facial image restoration by introducing Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that leverages user-specific photo galleries to improve denoising performance.
@@
 Our experiments, which included extensive evaluations on multiple challenging datasets, demonstrated that DiffPGD not only outperforms existing diffusion-based denoising methods but also delivers superior quantitative and qualitative results across various testing scenarios.
-This approach effectively mitigates false content generation in low-SNR conditions while bridging the gap in low-light imaging quality, ultimately paving the way for personalized and more robust image restoration techniques.
-Supplementary Material
-Figure 9: Simulation process for generating low-light noisy face images.
-Appendix A Dataset
-As mentioned in Section $4.1$ of the main paper, since no open-sourced low-light noisy datasets are available, we adopt a simulation approach to generate low-light noisy face images for training our model.
-Specifically, we utilize the CelebAMask-HQ dataset’s [30] face images as the ground truth RGB images.
-As shown in Fig.
-9, the first step in the simulation process involves passing the RGB image through a pre-trained Inverse ISP network [34] to generate a pseudo ground truth RAW mosaic image.
-For visualization purposes, we display the demosaiced version of this image.
-We utilize the pretrained weights provided by the authors [34], which were obtained by training the network on iPhone camera images, for unprocessing the RGB images into RAW format.
-This pseudo ground truth RAW image is then passed through a Poisson-Gaussian Simulator to simulate the low-light noisy RAW mosaic image.
-It is noteworthy that during the training process, we vary the camera sensor parameters (in the Poisson-Gaussian simulator) across a wide range of values to enable the model to handle diverse camera sensor noise profiles.
-The simulation setup assumes a $12$-bit image sensor with parameters similar to those of Canon DSLR cameras.
-The parameter ranges are summarized in Table 5.
-Table 5: Simulation parameters used in the Poisson-Gaussian Simulator.
-  Parameter Range
-Full Well Capacity $[19000\text{e}^{-},64000\text{e}^{-}]$
-Quantum Efficiency $[0.32,0.54]$
-Dark Current $[2.2\text{e}^{-},11.7\text{e}^{-}]$
-Read Noise $[2.2\text{e}^{-},10.8\text{e}^{-}]$
-Photons-per-Pixel (PPP) $[13\,\text{PPP},65\,\text{PPP}]$
-These parameters are carefully chosen to ensure realistic noise characteristics during training.
-Appendix B Additional Visual Results
-In Fig.
-12, Fig.
-14, and Fig.15, we present additional visual comparisons between our model and baseline methods on real-captured and simulated cases.
-Our results consistently demonstrate superior performance, particularly in cases of severe degradation.
-Figure 10: A network architecture diagram for identity feature restoration using a face encoder.
-  Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ID Score $\uparrow$
-w/ Encoder 12.62 0.2574 0.6315 0.6906
-w/ physical buffers 13.85 0.3753 0.6042 0.8039
-Table 6: Comparison of ID physical buffers and encoder-based methods on the real-captured photo test set.
-For a fair comparison, we use two same gallery images for each individual for both the encoder-based and physical buffer-based approaches.
-Appendix C More Implementation Details
-During training and testing, the image size is 256 $\times$ 256.
-The architecture of the diffusion model is adapted from [12].
-We modify the architecture so that the model can take the concatenation of physical buffers and degraded images as conditions.
-The model parameters are configured as follows: the number of channels is set to 128, channel multiple is (1, 1, 2, 2, 4, 4), head channels are 128, attention resolution is 16, dropout rate is 0.1, diffusion steps are 1000, and the two hyperparameters P2_gamma and p2_k are both set to 1.0.
-For additional architectural details, refer to [12].
-During inference, we adopt the DDIM method [46] with 100 denoising steps.
-The model is trained using the Adam optimizer [27] with a learning rate of $10^{-4}$.
-The weight decay is set to 0.
-All experiments are performed on a single NVIDIA A100 GPU.
-Appendix D Encoder vs. Physical Buffer?
-Many previous studies have used encoder-based methods to extract latent facial feature information from reference photos[51].
-To evaluate whether the encoder can effectively extract enough facial features to guide the diffusion model in solving the low-light enhancement and denoising task, we conducted an experiment.
-As illustrated in Fig.
-10, we utilized an encoder (ResNet-18 [19]) to extract latent codes from the target individual’s gallery photos.
-Following the method employed by ADM [12] for time embedding, we scaled and shifted the features in each layer of the diffusion model using the latent codes.
-This process integrates the identity information extracted by the encoder into the diffusion model.
-We compared the restored images generated with the encoder-based method with the proposed physical buffer-guided method on our real-captured photo set.
-This is a special semi-real set, in which we take photos of some printed paper photos at night.
-We prepare this semi-real set because The ground-truth will be available to be compared with.
-The quantitative and visual comparisons are presented in Table 6 and Fig.
-11, respectively.
-The results demonstrate that guiding the diffusion model using ID physical buffers produces significantly better identity preservation than the encoder-based approach.
-One major reason for this is that the ID physical buffers serve as robust conditions, enabling the extraction of comprehensive facial features from the gallery.
-In contrast, the encoder-based method is less focused on facial features and is often influenced by irrelevant factors such as background and hairstyle.
-Consequently, the latent codes extracted by the encoder are less effective in guiding the diffusion model to the personalized generative space compared to the ID physical buffers.
-Appendix E Gallery Photos
-Fig. 13 presents the the gallery photos for the test subjects included in the Fig. 6 of the main paper.
-Figure 11: Visual comparison on different identity representation strategies (physical buffer and face encoder).
-Figure 12: Additional real-captured visual comparison.
-Figure 13: Gallery photos.
-Figure 14: Additional visual comparison.
-Figure 15: Additional visual comparison.
+This approach effectively mitigates false content generation in low-SNR conditions while bridging the gap in low-light imaging quality, ultimately paving a way for personalized and more robust image restoration techniques.